### v3

# CoRe: Benchmarking LLMs’ Code Reasoning Capabilities through Static Analysis Tasks

Abstract
Large language models (LLMs) have been widely adopted across diverse domains of software engineering, such as code generation, program repair, and vulnerability detection. These applications require understanding beyond surface-level code patterns: value propagation, control flow, and interdependence between program elements. However, existing benchmarks primarily evaluate end-to-end outcomes, such as whether code is correctly repaired or generated, leaving the models’ ability for program semantic reasoning underexplored. This work presents CoRe, a high-quality, human-verified benchmark designed to evaluate LLMs on fundamental static analysis tasks. CoRe includes 12,553 task instances spanning data dependency, control dependency, and information flow across programs written in C/C++, Java, and Python. To ensure semantic diversity and reasoning complexity, we propose a semantics-aware diverse sampling strategy that selects targets and task instances based on structural coverage and dependency depth. We evaluate 10 mainstream LLMs and show that, while they perform well at identifying dependencies, models still struggle with tasks that require deeper semantic understanding and multi-step reasoning. We further conduct qualitative analyses to uncover key challenges, such as complex control structures and backward dependency patterns, offering insights into improving LLMs’ code reasoning capabilities.
1 Introduction
Large Language Models (LLMs) have shown remarkable capabilities across a wide range of domains (Achiam et al., 2023; Team et al., 2023; Wei et al., 2022; Yao et al., 2023; Grattafiori et al., 2024b; Xie et al., 2024, 2025). In the area of software engineering, they are increasingly used in tasks such as program synthesis Austin et al. (2021); Wu et al. (2024), program repair (Xia et al., 2024; Zhang et al., 2024c), and test generation (Xia et al., 2023; Schäfer et al., 2024). LLMs are usually prompted with high-level objectives Xia et al. (2024); Zhang et al. (2024c), such as identifying a buggy line, generating test cases to reach a target condition, or producing patches to fix faulty behavior.
Our Motivation of Benchmarking. Success in such tasks requires more than surface-level pattern recognition. Specifically, it requires a deep understanding of the program semantics: how values propagate through statements, how control structures govern program execution, and how different parts of the program influence one another. For instance, in the fuzzing scenario shown in Figure 1b, to trigger a vulnerability at line 12, the input must satisfy a nested series of conditions at lines 2, 4, 7, 9, and 10. Failing any of these checks leads to early termination or divergence from the path. This kind of control-dependent reasoning is essential for success in these tasks.
At the same time, researchers are beginning to use LLMs as static analyzers, applying them to downstream tasks such as vulnerability detection Nunez et al. (2024), automatic debugging Jiang et al. (2024a), and program repair (Li et al., 2025a). These approaches, no matter whether implicitly relying on or explicitly prompting LLMs for code reasoning, assume that the model has mastered core code semantic analysis skills, which can support downstream clients with exceptional performance empirically. However, their underlying ability to reason about program semantics remains under-evaluated despite promising end-task results. Therefore, we urgently require an effective benchmarking solution to directly assess whether LLMs possess the deep semantic understanding and reasoning capabilities necessary to support complex software engineering tasks.
Limitations of Existing Benchmarks. Existing benchmarks mainly target the evaluation of LLMs upon code-centric tasks in an end-to-end fashion (Chen et al., 2023; Ding et al., 2024a; Chen et al., 2021; Liang et al., 2025; Nguyen et al., 2025), without fundamentally assessing the model’s reasoning ability over program semantics. For example, program repair benchmarks such as SWE-Bench (Jimenez et al., 2024) and SWE-Lancer (Miserendino et al., 2025) provide buggy-fixed code pairs to evaluate patch generation correctness, yet lack fine‐grained ground-truth explanations of why each bug occurs. Other efforts that target fundamental program properties rather than downstream applications, like dynamic trace prediction (Yadavally et al., 2024a; Jiang et al., 2024b; Ding et al., 2024b), focus exclusively on observed runtime behaviors under specific inputs, leaving unexercised branches and other possible inputs overlooked. Consequently, they fail to provide adequate evaluation of whether LLMs possess the semantic reasoning required for advanced coding tasks, which highlights the need for a benchmark that directly evaluates core program analysis skills.
Our Benchmark. To fill this gap, we introduce CoRe, a high-quality, multi-lingual benchmark for evaluating LLMs’ Code Reasoning capabilities with fundamental static analysis tasks. CoRe includes human-verified task instances covering data dependency, control dependency, and information flow across C/C++, Java, and Python. The benchmark consists of 12,553 task instances drawn from 180 programs, selected through semantics-aware diverse sampling to ensure both task diversity and non-trivial reasoning complexity. Due to the inability of existing program analysis techniques to automatically and accurately extract comprehensive semantic properties from multi-languages, we adopt a semi-automated annotation pipeline supplemented with substantial human effort. This ensures the soundness and completeness of the annotations, resulting in a rigorously curated, high-quality dataset.
Empirical Results and Findings. Our extensive evaluation of 10 state-of-the-art LLMs, including 6 top reasoning models, along with qualitative analyses, reveals the key following:
  • Reasoning models consistently outperform non-reasoning models across most tasks, with a performance margin of 5.2–31.5%; Gemini 2.5 Pro achieves the best overall results.
  • Most of the models perform well on simpler tasks of identifying a dependency between two program elements, achieving F1 scores ranging from 68.9% to 92.56%.
  • Models struggle with tasks requiring deeper semantic understanding and multi-step reasoning, such as trace generation and source enumeration, with scores falling below 50%.
  • Performance drops significantly in the presence of complex control structures, longer function bodies, and backward or non-sequential dependency patterns, with a performance gap up to 49.5%.
  • When prompted to provide both classification and a trace, models are more likely to predict the existence of a dependency, indicating increased false positives under greater task complexity.
Availability. We release the dataset and code at https://corebench.github.io/ with Apache-2.0 license.
2 Background and Motivation
Figure 1: Real-world motivating examples for data dependency, control dependency, and information flow in security and software engineering applications.
Understanding program semantics requires reasoning beyond surface-level syntax. In traditional program analysis, dependencies are central to modeling how specific program values flow between different program lines, facilitating a wide range of applications, including bug detection(Zheng et al., 2024; Ayewah et al., 2008; Fan et al., 2019; Shi et al., 2023), program optimization (Chaitin, 1982; Cytron et al., 1991; Ahmed et al., 2024; Brandt and Schneider, 2009), and testing(Pacheco and Ernst, 2007; Böhme et al., 2017; Chen and Chen, 2018; Liu et al., 2024b, 2022; Fioraldi et al., 2020). In this work, we evaluate LLMs on three core reasoning tasks: data dependency, control dependency, and information flow. The first two represent fundamental program dependencies, while the third captures higher-level semantic reasoning about how program values propagate through execution. Evaluating LLMs on them provides insight into their ability to capture deep semantic structure in code.
Before introducing the definitions of the three dependencies and their applications, we define several basic units used throughout our benchmark.
Variables. A variable is denoted as $\texttt{name}_{line}$, where name is the identifier and line indicates the line number where the value of the variable is assigned or updated. For example, in the statement x = y + 1 on line 3, we refer to variable x as $\texttt{x}_{3}$, since it is assigned a new value.
Line Numbers. We use $\ell_{i}$ to refer to the $i$-th line of code. This notation is used when describing relationships between statements based on their locations, particularly in control dependency.
Variable Relationships. We use $\rightarrow$ to denote a direct relationship between two variables or statements, including data dependencies ($\rightarrow_{D}$), control dependencies ($\rightarrow_{C}$), and information flows ($\rightarrow_{I}$). A data dependency trace is written as a sequence such as $\texttt{x}_{1}\rightarrow_{D}\texttt{y}_{3}\rightarrow_{D}\texttt{z}_{5}$, where each arrow indicates that the target variable is directly data-dependent on the source. We use $\rightsquigarrow$ to denote a transitive relationship, indicating that one variable is indirectly influenced by another through one or more intermediate steps. For example, the above trace implies $\texttt{x}_{1}\rightsquigarrow_{D}\texttt{z}_{5}$.
2.1 Data Dependency
A data dependency occurs when the value of one variable depends on the value of another, typically arising when variables are assigned and then subsequently used Alfred et al. (2007).
1x = 5
2y = x + 2
3z = y + 3 # z depends on y, and transitively on x
Here, $\texttt{y}_{2}$ is directly data dependent on $\texttt{x}_{1}$, and $\texttt{z}_{3}$ is directly data dependent on $\texttt{y}_{2}$. This forms a transitive data dependency trace: $\texttt{x}_{1}\rightarrow_{D}\texttt{y}_{2}\rightarrow_{D}\texttt{z}_{3}$. Therefore, $\texttt{z}_{3}$ has an indirect data dependence on $\texttt{x}_{1}$, denoted as $\texttt{x}_{1}\rightsquigarrow_{D}\texttt{z}_{3}$.
While this example illustrates a simple chain of assignments, real-world programs often exhibit more complex forms of data dependence involving pointers, aliasing, and indirect memory access. Accurately capturing such dependencies remains a key challenge for both traditional static analysis tools and LLM-based reasoning systems.
Applications. Data dependencies essentially induce the def-use chains in the program, serving as the ingredient for diverse static analysis applications, such as program slicing (Yadavally et al., 2024a), compiler optimizations (Ferrante et al., 1987), and taint tracking in security (Chen and Chen, 2018; Arzt et al., 2014), with growing interest in using LLMs to assist such analyses (Wang et al., 2024a; Li et al., 2025b).
Fig. 1a presents an example of a real-world security vulnerability CVE-2017-5638 in Apache Struts2, which is caused by the usage of a program value that is data-dependent on an untrusted input. Here, $\texttt{header}_{2}$ receives untrusted input from an HTTP request and is used to define $\texttt{expr}_{4}$, forming a direct data dependency: $\texttt{header}_{2}\rightarrow_{D}\texttt{expr}_{4}$. The value of $\texttt{expr}_{4}$ is then used at $\ell_{6}$ in a critical operation. This example demonstrates how data dependency analysis supports taint analysis by uncovering how untrusted input propagates through variable assignments to reach sensitive operations, making it a key technique for detecting and analyzing security vulnerabilities.
2.2 Control Dependency
Control dependency captures whether the execution of one statement is governed by another (Podgurski and Clarke, 1990). We use the line numbers of the statements to indicate the dependency. Specifically, a statement at $\ell_{2}$ is control-dependent on $\ell_{1}$ if the condition at $\ell_{1}$ determines whether $\ell_{2}$ executes. This holds when there is at least one branch from $\ell_{1}$ where $\ell_{2}$ always executes, and another where it may not.
1if x > 0: # controls execution of line 2
2 if y > 0: # controls execution of line 3
3 a = 3 # only executes if both conditions hold
4b = a + 1 # executes unconditionally
Here, $\ell_{3}$ is directly control dependent on $\ell_{2}$, and $\ell_{2}$ on $\ell_{1}$, forming the trace $\ell_{1}$ $\rightarrow_{C}$ $\ell_{2}$ $\rightarrow_{C}$ $\ell_{3}$. This trace represents a transitive control dependency: $\ell_{1}$ $\rightsquigarrow_{C}$ $\ell_{3}$. $\ell_{4}$ is not control dependent on any condition in this code snippet. In our benchmark, tasks targeting control dependencies require LLMs to reason about how branching conditions affect execution paths. This involves identifying statements whose execution depends on earlier control conditions, including transitive cases across nested or compound conditionals.
Applications. Control dependencies capture the behavior of branches. They are widely used in path reachability reasoning in various program analysis tasks, such as bug detection Chen et al. (2019); Zhang et al. (2023) and program verification Jensen et al. (1999); Chen and He (2018), which involves reasoning about the variables affecting the path execution.
Fig. 1b is an example of applying control dependency analysis in fuzzing, which presents a real-world vulnerability CVE-2022-26129 in the BABEL routing daemon. The call to memcpy at $\ell_{12}$ is the vulnerable operation that may cause a buffer overread. Reaching this line requires the input to satisfy a chain of conditions from $\ell_{2}$, $\ell_{4}$, $\ell_{7}$, $\ell_{9}$, and $\ell_{10}$. These conditionals form a transitive control dependency trace ending at $\ell_{12}$. Because triggering the vulnerable memory access depends on passing multiple branching checks, control dependency analysis is crucial for guiding fuzzers to explore such paths and uncover deep bugs Huang et al. (2020); Jiang et al. (2023).
2.3 Information Flow
Information flow (Denning and Denning, 1977; Genaim and Spoto, 2005) captures how the value of one variable can influence another through data or control dependencies. It may be explicit, as in direct assignments, or implicit, when control conditions determine which value a variable receives. An information flow may be a sequence of implicit, explicit, or mixed flows.
1x = 0
2if x == 0:
3 y = 0 # y is assigned if x is true, therefore a implicit flow from x
4else:
5 y = 1 # alternative assignment to y
6z = y + 1 # explicit flow from y
Here, the condition at $\ell_{2}$ directly reads $\texttt{x}_{1}$ and has direct control dependence over both $\texttt{y}_{3}$ and $\texttt{y}_{5}$, so $\texttt{x}_{1}\rightarrow_{I}\texttt{y}_{3}$ (or $\texttt{y}_{5}$) is an implicit flow. $\texttt{y}_{3}\rightarrow_{I}\texttt{z}_{6}$ (or $\texttt{y}_{5}\rightarrow_{I}\texttt{z}_{6}$) is an explicit flow through assignment. Combining both gives a transitive information flow: $\texttt{x}_{1}\rightsquigarrow_{I}\texttt{z}_{6}$. In our benchmark, tasks targeting information flow require LLMs to reason over both explicit assignments and implicit control-induced flows, and to recover transitive chains of influence between variable definitions.
Applications. Information flow analysis is widely used in security and privacy, including taint tracking, non-interference, and enforcement of confidentiality and integrity policies Sabelfeld and Myers (2003). In software engineering, it supports the program slicing and more downstream tasks, such as program debugging and fault localization. For example, it helps isolate the minimal subprogram relevant to a fault, improving code understanding and the developer’s ability to identify and fix bugs (Francel and Rugaber, 2001; Yadavally et al., 2024b). Recent research increasingly explores LLMs for assisting such analyses (Mashnoor et al., 2025).
Fig. 1c is an example of slicing for debugging and fault localization, based on a real-world code snippet [^1] Originally from StackOverflow post #16180130. : At $\ell_{13}$, the variable in may be null, leading to a potential NullPointerException (NPE). A static backward slice rooted at $\ell_{13}$ with respect to in reveals two possible assignments: $\ell_{2}$ and $\ell_{10}$. Whether these assignments execute depends on the condition at $\ell_{1}$, which govern the control flow paths. Thus, the slice includes both the relevant data dependencies and the control dependencies that determine whether in is properly initialized.
3 Our Benchmark: CoRe
We introduce a human-verified high-quality benchmark for evaluating the semantic reasoning capabilities of LLMs in code. Unlike prior benchmarks targeting code generation or functional correctness, our objective is to provide a more fine-grained evaluation of LLMs’ code semantic reasoning ability by investigating the three core semantic properties: data dependency, control dependency, and information flow. The benchmark is multi-lingual (C/C++, Java, and Python), and diverse in both task type and difficulty. It contains 12,553 tasks derived from 180 annotated programs, each manually curated or reviewed. A semantics-aware sampling strategy guides the construction of the task, ensuring both structural diversity and reasoning complexity. By emphasizing reasoning over program structure and behavior, this benchmark provides a targeted testbed to evaluate LLMs’ understanding of the code semantics.
3.1 Benchmark Construction
We sample programs from two major sources: CodeNet (Puri et al., 2021), a large-scale dataset of competitive code, and Google Code Jam (GCJ) (40), which contains expert-written solutions to algorithmic problems. From these, we select 60 files each in C/C++, Java, and Python, resulting in 180 functions in total. We also select one target function from each file.
Our sampling ensures diverse program structures and a balanced distribution of lines of code (LoC) in each programming language. Specifically, we categorize candidate functions into four LoC buckets: 21–40, 41–60, 61–80, and 81–100, and ensure equal distribution among them. As function calls can introduce long dependency chains across functions, making it difficult to provide all the relevant functions in a single prompt, we only focus on intra-procedural analysis, excluding inter-procedural reasoning. We only select the functions that do not invoke any non-library functions, which is also a common practice in existing studies (Yadavally et al., 2024a, b), Check appendix A for benchmark construction details.
3.2 Data Annotation
For each function, we select a target variable and label its direct data and control dependencies. All variables transitively involved in these dependencies are also annotated. Due to the infeasibility of automating annotation (Appendix B.1), we adopt a semi-automated approach. A custom static analysis tool built on tree-sitter generates initial labels, which are then manually validated by two authors, each with over 5 years of program analysis experience. Conflicts are resolved by a third author. This process achieves an agreement rate of 87.5 %. In total, it yields 6,306 annotated variables and 48,050 lines of annotation. More annotation details are provided in the Appendix B.
3.3 Task Design
We define three task types corresponding to core static analysis concepts: data dependency, control dependency, and information flow (Section 2). Each task type consists of multiple task instances. A task instance is a query that asks the model to analyze a specific type of dependency between program elements within a given program.
3.3.1 Task Formulation
While our raw annotations can support various static analysis tasks, we define and focus on two query types for each dependency category. Exact question formulations are provided in Appendix C.
Pairwise Query. Given two program elements (variables or lines), determine whether a specific dependency exists. If so, the model must also output a valid trace: a transitive sequence of direct dependencies or flows from source to target. When multiple traces exist, any one is sufficient.
Target-Centric Query. Given a single program element, list all other elements in the same function that have the specified dependency relation over it (e.g., all variables that the target is data dependent on). The model is expected to return a complete (orderless) set.
3.3.2 Task Instance Generation with Semantics-Aware Diverse Sampling
Table 1: CoRe task instance distribution
Task Type Postive Negative Total
Data Dependency 1,814 2,800 4,614
Control Dependency 1,693 1,959 3,652
Information Flow 2,291 1,996 4,287
Total 5,798 6,755 12,553
Table 2: CoRe Lite task instance distribution
Task Type Postive Negative Total
Data Dependency 209 381 590
Control Dependency 239 250 489
Information Flow 291 214 505
Total 739 845 1,584
For each of the 180 annotated programs, we generate task instances for all three task types. For each task type, we sample up to five targets per program: variables for data dependency and information flow, and lines for control dependency. For each target, we construct up to five positive and five negative task instances. Each task instance poses a query about whether another variable or line in the same program holds a specific dependency relationship over the target.
We apply a Semantics-Aware Diverse Sampling strategy to guide both target and instance selection. The strategy is informed by code structure, such as control flow, as well as semantic dependencies to ensure both diversity and reasoning complexity. It promotes diversity by avoiding overlapping traces or repeated contexts, and enforces reasoning complexity by selecting targets with non-trivial dependency structures and negative instances that are structurally plausible to test the model’s ability to distinguish true dependencies from misleading patterns.
This process results in a total of 12,553 task instances. Table 2 summarizes the distribution across task types and programming languages. Ground truth labels are automatically derived from the annotation. See Appendix D for sampling algorithms, statistics, and ablation study.
3.4 CoRe Lite
To encourage broader adoption of CoRe, we release a lite subset containing 1,584 task instances (Table 2). Based on the full benchmark constructed in Section 3.3.2, we randomly sample one instance per target, providing a smaller yet diverse and representative version. We evaluate the target-centric query only on CoRe Lite, since each target appears only once, avoiding redundancy.
4 Experimental Setup
Prompt Design. Each prompt includes detailed definitions, the expected output format, and 5–7 illustrative examples drawn from small synthetic programs designed for demonstration (Appendix J). Also, each example includes step-by-step explanations and final outputs.
Table 3: Evaluated models with sizes and reasoning capabilities.
Claude 3.7 Claude 3.5 DeepSeek R1 DeepSeek V3 Gemini 2.5 Pro GPT o3 GPT o4-mini GPT 4o Llama3.1 Qwen3
Size - - 671B 671B - - - - 405B 235B
Reason
Models. Table 3 shows the 10 models we evaluate in this paper. Given the complexity and reasoning-oriented nature of our benchmark, we primarily focus on models with reasoning capabilities and strong coding performance, typically with sizes above 200B parameters. In total, we evaluate 6 reasoning and 4 non-reasoning models. Some non-reasoning models still show strong code performance despite lacking explicit reasoning abilities. Invocation and API details are in Appendix E.
Evaluation Metrics. We evaluate model performance across three aspects: dependency classification, trace quality, and dependency source enumeration.
  • Dependency Classification. Model is queried whether a specific dependency relation holds between two program elements (variables or lines). We report precision, recall, and F1 score.
  • Trace Quality. If a dependency is predicted to exist, the model must also produce a valid trace: a transitive sequence of direct dependencies or flows. We evaluate this using Correct Trace Rate (CT), defined as the fraction of traces that are entirely correct.
  • Dependency Source Enumeration. Given a target variable or line, the model is asked to list all elements (variables or lines) that influence it via a specific dependency type. We report Exact Match (EM): the percentage of predictions that match the full ground-truth set.
Fine-grained metrics such as edge-level correctness and partial set recall are defined in Appendix F, where we also report complete results on those to support deeper analysis.
5 Evalution Results
We organize our evaluation around three questions: RQ1 (Section 5.1) reports overall performance, RQ2 (Section 5.2) presents a qualitative study of factors affecting model behavior, and RQ3 (Section 5.3) examines the impact of different experimental designs.
5.1 RQ1: How Well Do LLMs Perform on Dependency Reasoning Tasks?
Table 4: Model performance on various tasks on three task types: Data Dependency (Data), Control Dependency (Control), and Information Flow (InfoFlow) on CoRe Lite.
Models
(Reasoning v.s
Dependency Classification
F1 (%)
Trace Generation
Correct Trace Rate (%)
Dependency Source Enumeration
Exact Match (%)
Non-Reasoning) Data Control InfoFlow Overall Data Control InfoFlow Overall Data Control InfoFlow Overall
Claude 3.7 76.94 88.84 81.57 82.07 69.29 60.04 39.90 57.13 20.58 51.70 6.92 25.82
DeepSeek R1 83.29 92.28 83.59 86.18 67.31 66.62 39.58 58.36 38.88 48.37 7.12 31.82
Gemini 2.5 Pro 88.53 92.49 94.79 91.74 90.38 92.26 68.66 84.02 49.43 75.66 26.73 50.25
GPT o3 93.24 92.11 92.13 92.56 86.23 77.52 52.37 72.80 41.89 70.90 15.61 42.61
GPT o4-mini 84.70 91.98 84.08 86.74 70.39 66.83 42.32 60.43 29.11 61.76 8.90 32.89
Qwen3 235B 82.33 89.00 69.51 80.31 65.48 59.19 29.85 52.26 21.36 44.61 4.74 23.30
Claude 3.5 70.67 78.36 84.06 77.27 54.19 57.34 35.90 49.43 11.14 40.46 7.72 19.13
DeepSeek V3 71.77 79.48 76.88 75.80 56.37 40.27 23.66 41.08 15.95 21.33 2.58 13.38
GPT 4o 69.75 77.11 76.63 74.16 61.85 43.15 22.53 43.56 12.28 29.06 2.97 14.52
Llama 3.1 405B 63.15 70.30 74.49 68.93 39.50 27.77 15.98 28.48 2.96 5.35 1.59 3.28
(a) Dependency Classification
(b) Trace Generation
(c) Dependency Enumeration
Figure 2: Performance distribution of reasoning vs. non-reasoning models across all three task types: Data Dependency (Data), Control Dependency (Control), and Information Flow (InfoFlow).
Tables 4 report the performance of the 10 evaluated models on CoRe Lite, split into reasoning (top) and non-reasoning (bottom) models. We evaluate dependency classification (whether a dependency exists), trace generation (producing a valid transitive trace), and dependency source enumeration (listing all sources for a given target) across data dependency, control dependency, and information flow. Full results on CoRe with all metrics and more analysis are in Appendix G.
Reasoning models generally perform well on dependency classification, achieving over 80% F1 score. However, trace generation remains challenging: only Gemini 2.5 Pro and GPT-o3 achieve over 70% correct rate, while others lag at 20–60%. Dependency source enumeration appears to be the hardest, with most reasoning models performing below 40%, and non-reasoning models below 20%. Among all models, Gemini 2.5 Pro shows the best overall performance, followed by GPT-o3.
Fig. 2 shows the performance distribution across all tasks. Reasoning models consistently outperform non-reasoning ones by a margin of 5.2–31.5%. Despite achieving strong classification F1 for information flow (Fig. 2(a)), models often struggle to produce correct traces for it (Fig. 2(b)), which requires integrating both explicit and implicit flows. A similar gap appears for control dependency, likely resulting from challenges in parsing control structures, especially under early returns or nested blocks (Section 5.2). Control dependency enumeration (Fig. 2(c)) is relatively easier due to its limited set of line-level sources (avg. 4.0), while data and information flow require enumerating more variable-level sources (avg. 7.7, 17.0), making them much more difficult.
As shown by the above statistics, while reasoning models can identify dependencies with high precision, generating accurate traces and enumerating all relevant sources remains a significant challenge, especially for information flow, which is the most challenging task as it requires reasoning over both data propagation and control influences.
5.2 RQ2: What Factors Influence Model Performance?
(a) Function Length
(b) Number of Conditions
(c) Early Exit
(d) Reverse Dep.
Figure 3: Factors affect model performance, with longer functions, complex control, and reverse dependencies leading to consistent drops.
To better understand the challenges of the benchmark, we analyze factors that increase task difficulty and assess their impact on model performance.
Function Length.
As shown in Fig. 3(a), model performance (correct trace rate) on trace generation consistently declines as the length of the function increases. The performance gap between short (21–40 LoC) and long (81–100 LoC) functions reaches 8.9–15.8%. Longer functions present more tokens for the model to process, requiring it to reason over larger and often more entangled code contexts, including deeply nested logic and scattered dependencies.
Control Structure Complexity. We study how control structures affect trace generation quality. Fig. 3(b) shows the relationship between the number of conditions (e.g., if, while, for) in a function and the correct trace rate. As the number of conditions increases, model performance drops consistently, with differences of 24.4–44.6%.
We also examine the impact of early exits such as return or break. In Fig. 3(c), we compare the exact match rate of dependency source enumeration between cases where the target variable is influenced by early exit logic and those where it is not. Across all models, we observe a consistent drop in performance of up to 21.7%.
These findings suggest that complex or non-linear control flows significantly increase reasoning difficulty, likely due to the need to track multiple execution paths and conditionally gated behaviors.
Reverse Dependency Order. With the pair-wise query, each query involves two program elements (variables or lines), such as “Does variable $a$ have a data dependency over variable bb?” Typically, the source element ($a$) appears earlier in the code than the target ($b$), matching natural expectations (i.e., line($a$) << line($b$)). However, reverse dependencies also occur, especially in loops, where a later variable can influence an earlier one. Fig. 3(d) shows how this affects classification F1 scores. When the source appears after the target, model performance drops significantly, by up to 49.5% (Claude 3.7). This suggests that LLMs are biased toward left-to-right reasoning and struggle with abstract dependency graphs that require backward or bidirectional analysis.
5.3 RQ3: How Do Different Experimental Settings Affect Model Behavior?
We evaluate the impact of three experimental variations on model performance, focusing on how task framing and input context influence model behavior.
Asking for Traces vs. Classification Only. In the default setting with pair-wise query, models are asked to predict whether a dependency exists and provide a supporting trace. We compare this to a simplified classification-only setup, where no trace is requested. As shown in Fig. 4(a), this results in higher precision but lower recall. This suggests that when prompted to produce a trace, models are more likely to answer “yes” (i.e., assert that the relationship holds).
Clipping the Function Context. By default, the entire file is provided as input, and the prompt specifies the target function. We compare this to a clipped setup that includes only the target function. Fig. 4(b) shows that models perform consistently better in the clipped setting, improving by up to 5.7%, indicating that models struggle to focus when given longer irrelevant context, even with a specified function, in contrast to traditional analysis tools, which inherently isolate scope.
Few-Shot Learning (FSL). While our base prompt includes examples and definitions, we further explore an FSL setting using additional examples retrieved via sparse and dense retrievers. Surprisingly, results do not improve and even degrade, likely due to (1) longer inputs distracting from core instructions, and (2) potential bias from retrieved examples, even with balanced labels. Additionally, standard retrievers fail to capture structural or semantic similarity between tasks, suggesting that better structure-aware retrieval is needed. Details are presented in Appendix 4.
6 Related Work
6.1 LLMs for Software Engineering Tasks
LLMs have been widely applied across many domains Brown et al. (2020); Zhang et al. (2024b); Wang et al. (2024b); Beurer-Kellner et al. (2023), including software engineering, where they are used for tasks such as programrepair (Xia et al., 2024; Zhang et al., 2024c; Xia et al., 2023; Zhang et al., 2024a), code generation Jiang et al. (2024a); Liang et al. (2025); Lyu et al. (2024); Pei et al. (2024); Gao et al. (2025), software testing (Schäfer et al., 2024; Yang et al., 2024; Zheng et al., 2025b, c, a), and vulnerability detection (Chen et al., 2023; Li et al., 2025b; Du et al., 2024). A key enabler for these applications is reasoning over static program properties. Traditional static analysis tools, however, are language-specific and require extensive manual engineering. To overcome these limitations, recent works increasingly use LLMs themselves to perform static analysis as an intermediate step. LLMDFA (Wang et al., 2024a) and E&V (Hao et al., 2023) apply LLM-based data- and control-flow reasoning to detect bugs; IRIS Li et al. (2025b) leverages LLMs to examine data-flow paths reported by static analyzers, effectively mitigating the false positives. CoRe offers a fine-grained benchmark to evaluate LLMs’ capabilities in reasoning about program semantics, which underlie these downstream SE tasks.
6.2 Code Reasoning Benchmark and Evaluation
Code reasoning, the ability to understand program behavior, is fundamental to downstream LLM applications such as program synthesis (Huang et al., 2024; Bhaskar et al., 2023; Du et al., 2023), program repair (Jimenez et al., 2024; Ouyang et al., 2024; Wu et al., 2023), and bug or vulnerability detection (Yildiz et al., 2025; Tamberg and Bahsi, 2025; Chen et al., 2023; Ding et al., 2024a). These tasks require models to reason over program semantics, including data and control flow. Yet, most benchmarks only measure downstream success (e.g., detecting vulnerabilities successfully) without directly probing models’ reasoning capabilities on code semantics.
Recent work has begun to explicitly evaluate internal code reasoning. CRUXEval Xu et al. (2025) focuses solely on Python and evaluates input-output behavior, which aligns more with dynamic analysis. In contrast, our benchmark targets static analysis skills and covers Python, Java, and C/C++. CRQBench (Dinella et al., 2024) evaluates LLMs’ semantic reasoning with C++ questions focused on variable tracking and behavioral equivalence. However, its questions focus on high-level behaviors and are automatically generated from pull requests and review comments, making them error-prone. Similarly, (Pei et al., 2024) evaluates finetuned LLMs on several program analysis tasks, such as memory region prediction and function signature recovery from Java programs. However, the tasks remain high-level. CodeMMLU Nguyen et al. (2025) focuses on high level also focuses on knowledge-based high-level questions. REval Chen et al. (2025) focuses on programs’ runtime behaviors such as code coverage prediction. CodeCrash (Lam et al., 2025) evaluates model robustness via adversarial code perturbations, but does not directly assess semantic reasoning. CoRe fills this gap by enabling fine-grained assessment of models’ fundamental code understanding capabilities, including data and control dependencies and information flow, beyond surface-level pattern recognition or input-output matching.
7 Conclusion
We introduce CoRe, a high-quality multilingual benchmark for evaluating LLMs on fundamental static analysis tasks, including data dependency, control dependency, and information flow across C/C++, Java, and Python. It includes 12,553 human-verified, diverse instances from 180 programs. Our evaluation of 10 state-of-the-art LLMs shows that while many models perform well at identifying dependencies, they struggle with deeper semantic tasks such as generating valid traces. Our qualitative analysis further highlights key challenges such as long functions, complex control structures, and backward reasoning. CoRe fills an important gap by enabling fine-grained assessment of fundamental program reasoning skills beyond surface-level cues.
Future Directions.
A natural extension is to support inter-procedural analysis. Despite the difficulty of obtaining reliable annotations, one possible approach is to employ lightweight static analysis tools to identify function pairs with dependencies (e.g., via calling context). Once each function is fully annotated, a call graph can be constructed to enable inter-procedural reasoning.
Another promising direction is to examine the resilience of LLMs’ code reasoning under various forms of code obfuscation or subtle adversarial perturbations. Such analysis could shed light on the robustness and reliability of their semantic understanding in complex real-world scenarios.
8 Limitations
Our benchmark focuses on relatively short (within 100 LoC) due to the difficulty of scaling automated construction (Section B) and significant manual effort in annotation (Section 3.2). However, the functions within this size range still reflect real-world practices, as well-structured codebases often keep functions concise for maintainability (e.g.,  40–50 LoC) Nvidia (2025); Google (2025). Nonetheless, we curate 12,553 instances with an assurance of diversity and difficulty (Section 3.3.2). While annotation is manual and may introduce errors, we alleviate this through sufficient cross-validation (Section 3.2).
9 Acknowledgement
We are grateful to the Center for AI Safety for providing computational resources. This work was funded in part by the National Science Foundation (NSF) Awards 2006688, 1901242, SHF-1901242, SHF-1910300, Proto-OKN 2333736, IIS-2416835, DARPA VSPELLS - HR001120S0058, ONR N00014-23-1-2081, and Amazon. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
Appendix A Benchmark Construction
A.1 Dataset Sources
CodeNet.
CodeNet [73] is a large-scale dataset introduced by IBM. It consists of over 14 million code submissions spanning 55 programming languages, with a total of approximately 500 million lines of code. The dataset is derived from competitive programming problems and includes rich annotations such as problem IDs, programming language labels, code correctness (based on test case results), runtime and memory usage, and input-output test cases for the majority of problems. These annotations make CodeNet particularly valuable for supervised learning tasks and for constructing benchmarks focused on correctness, efficiency, or cross-language analysis. In our work, we use CodeNet as a source of diverse, well-labeled, real-world functions that reflect authentic coding practices across a variety of problem types and difficulty levels. The dataset is licensed under Apache 2.0.
Google Code Jam (GCJ).
Google Code Jam [40] is a long-running global programming competition hosted by Google, which features algorithmic problems of varying complexity. GCJ problems are typically accompanied by official test sets and are designed to test a range of skills including algorithm design, edge-case handling, and optimization under constraints. Solutions submitted by participants are generally concise, high-quality, and written under time pressure, providing a rich source of clean, focused code samples. For our benchmark, we extract complete solution files from GCJ archives, which offer naturally diverse control and data flow patterns that are especially well-suited for program analysis tasks. Specifically, we only sample the code from 2018 – 2021 collected from a public GitHub repository  [^2] https://github.com/Jur1cek/gcj-dataset that crawled data from the GCJ website. The repository is with an MIT license.
A.2 Data Sampling
Table 5 shows the distribution of sampled functions by dataset, language, and function length. We sample 180 program files in total, 90 from CodeNet and 90 from Google Code Jam (GCJ), and select one target function from each file for annotation.
Among the remaining files, we first remove all the comments and empty lines, and categorize candidate functions into four LoC buckets: 21–40, 41–60, 61–80, and 81–100. We enforce an even distribution across all buckets and across the three target languages (C/C++, Java, Python), resulting in a balanced and diverse set of functions suitable for dependency analysis.
To ensure consistent analysis scope, we discard samples that involve inter-procedural constructs such as function calls with non-local context or reliance on global variables. We also exclude files longer than 300 lines of code to maintain manageable complexity.
When necessary, we apply minimal, semantics-preserving edits to facilitate annotation. For example, we expand inline expressions or break compound statements (e.g., multiple statements on the same line separated by “;” in C or Java) into separate lines. These edits help standardize the structure of the code while preserving its original behavior.
Table 5: Distribution of sampled functions by dataset, language, and function length (in LOC).
Dataset Source Lang. Target Function LoC Total
21–40 41–60 61–80 81–100
CodeNet C/C++ 7 7 6 10 30
Java 8 7 8 7 30
Python 7 6 7 10 30
Google Code Jam (GCJ) C/C++ 8 8 9 5 30
Java 7 8 7 8 30
Python 8 9 8 5 30
Total 30 30 30 30 180
Appendix B Data Annotation
B.1 Challenges for Automation
As discussed in Section 2, the program properties addressed by our reasoning tasks are fundamentally semantic properties. According to Rice’s Theorem [74], determining whether a semantic property holds or not is inherently undecidable. Consequently, it is infeasible to develop an algorithm that fully automates the data annotation process and yields the prefect ground truth of the targeted semantic properties. Initially, we also explored employing static and dynamic analysis techniques to automatically generate data and control dependencies. However, we discovered that existing techniques are insufficient to significantly alleviate the annotation overhead.
Existing Static Analyzers.
Existing static analyzers, such as SVF [80], FlowDroid [6], CodeQL [8], and KLEE [16], predominantly analyze intermediate representation (IR) code rather than source code directly. IR code, which is generated by compiler frameworks, facilitates foundational analyses like pointer analysis [79]. Hence, the outputs of these analyzers typically depict the relations among program values within IR code rather than the values of source-level variables and expressions. However, LLMs operate directly on source code, not IRs, creating a semantic gap between traditional analysis tools and the kind of source-level reasoning we aim to evaluate. As a result, existing static analysis tools cannot be directly used to generate the labels required for our benchmark.
Furthermore, existing static analyzers are primarily designed for languages like C/C++ and Java, leaving Python, widely prevalent in the machine learning community, largely unaddressed. To the best of our knowledge, no practical Python static analyzer can accurately derive the specific semantic properties targeted by this work. Hence, utilizing static analyzers to automate data annotation is currently impractical.
Existing Dynamic Analyzers.
Dynamic analysis [32, 10, 78], unlike static analysis, is generally more language-agnostic. For given inputs, it readily tracks variable and expression values throughout program execution as long as we have the execution environment. Nevertheless, dynamic analysis only captures program behavior corresponding to specific inputs, potentially leaving many program behaviors unexplored. As underscored by prior dynamic testing research [46, 108], achieving sufficient coverage to comprehensively characterize all program behaviors is challenging. Despite advancements such as fuzz testing aimed at increasing coverage [109], complex control-flow structures, such as branches and loops, still significantly restrict the completeness of collected data and control dependencies.
Considering these limitations, we abstain from employing conventional static or dynamic analysis techniques for automating data annotation. Instead, we introduce the methodology detailed in Section 3, designed specifically to generate diverse, high-quality datasets applicable across multiple programming languages.
B.2 Annotation Rules
For each annotated data point (i.e., a sampled function), we randomly select one target variable. We then annotate all its direct data and control dependencies, and recursively expand the annotation to include all transitive dependencies. This ensures that complete dependency traces can be constructed during evaluation. We only annotate data and control dependencies, as information flow can later be derived from these annotations. The annotation strictly follows the definitions introduced in Section  2.1 and 2.2.
An example of a sampled C function from CodeNet  [^3] The sampled function main, spanning lines 12–40, is from CodeNet problem p00496, solution s700056700. and its corresponding annotation is shown below.
10...
11
12int main()
13{
14 int a[3001];
15 int b[3001];
16 int n, t, s, i, j, k, ans;
17 fgets(p=buf, 30, stdin);
18 n = getint();
19 p++;
20 t = getint();
21 p++;
22 s = getint();
23 for (i = 1; i <= n; i++) {
24 fgets(p=buf, 30, stdin);
25 a[i] = getint();
26 p++;
27 b[i] = getint();
28 }
29 ans = 0;
30 for (i = 1; i <= n; i++)
31 for (j = 1; j <= t; j++) {
32 dp[i][j] = max(dp[i-1][j], dp[i][j-1]);
33 k = j-b[i]; // <- The target variable
34 if (k >= 0 && (s <= k $|$ j <= s))
35 dp[i][j] = max(dp[i][j], dp[i-1][k] + a[i]);
36 ans = max(ans, dp[i][j]);
37 }
[…]
39 return 0;
40}
[…]
$\begin{lstlisting}$[language=Python, numbers=left, firstnumber=1]
target_var: k 33
other_vars:
- a 25
- p 26
- dp 32
- dp 35
variables:
k 33:
datadep:
- j 31
- b 27
- i 30
controldep:
- j 31
- t 31
b 27:
datadep:
- i 23
controldep:
- i 23
- n 23
t 31:
reset: False
datadep:
- t 20
controldep:
- i 30
- n 30
- j 31
- t 31
We annotate control dependencies at the variable level rather than the line level, as variable-level granularity is necessary for deriving information flow. Variables are represented in the (name, line) format (similar to the format introduced in Section 2), referring to where the variable is (re)defined or updated. We do not annotate variables that are only read at a line (e.g., y used in x = y + 1 without being updated).
Annotations are stored in a human-readable YAML format to facilitate manual editing and automated parsing. The YAML files consist of:
  • target_var: the selected target variable. Here it is (k,33), which is variable k defined in line 33 in the source code.
  • other_vars: variables annotated as irrelevant. In addition to annotating all variables involved in direct or transitive dependencies, annotators also labeled a set of irrelevant variables to serve as negative examples. These include variables that appear before the target line or within the same loop but are NOT semantically dependent on the target. These serve as negative examples to test a model’s ability to distinguish true semantic influence from structural proximity.
    For example, the variable a defined at line 25, (a,25), has no data or control dependencies leading to the target variable (k,33). That is, there is no information flow from (a,25) to (k,33). As a result, (a,25) is labeled as an other_var, as shown on line 3 in the annotation.
  • variables: each variable annotated with its direct data dependencies (datadep) and control dependencies (controldep).
Note that while we only annotate variables at the lines where they are (re)defined, an exception arises when assigning control dependencies. Specifically, the control-dependent lines for a variable may not themselves redefine that variable. For example, variable (k,33) is directly control-dependent on line 31. Therefore, it is annotated with control dependencies (j,31) and (t,31), even though t is not assigned or updated at line 31. To represent this dependency consistently using our (name, line) format, we include such cases in the annotation but mark them with “reset:False” (line 23 in the annotation). This flag ensures that these entries are excluded from data dependency processing while still being used for control and information flow reasoning.
B.3 Annotation Statistics
On average, each function contains 35.03 annotated variables as participating in a dependency or information flow relation. Each annotated variable has 1.92 direct data dependencies and 1.92 direct control dependencies.
The average manual annotation time is 37 minutes per annotator per function. To ensure annotation quality, each data point is cross-verified by at least two authors. Disagreements are resolved by a third author, with an observed agreement rate of 87.5 %.
Appendix C Task Formulation
We list the query templates used for both pairwise dependency and dependency source enumeration queries across data dependency, control dependency, and information flow, with the task-specific content (e.g., code snippets, variable names) highlighted in blue.
Note that these are only the query templates. For the full prompt templates used in our experiments, including task definitions, instructions, and examples, please refer to Appendix J.
Query for Pairwise Dependency Query (Data Dependency) Below is your target code snippet.
<target code with line number>
Question: Does (src_var, src_line) have data dependence over (dst_var, dst_line) in function target_function_name? If so, provide a trace.
Output:
Query for Dependency Source Enumeration (Data Dependency) Below is your target code snippet.
<target code with line number>
Question: Which variable instances have data dependence over (dst_var, dst_line) in function target_function_name? List all such variables.
Output:
Query for Pairwise Dependency Query (Control Dependency) Below is your target code snippet.
<target code with line number>
Question: Does src_line have control dependence over dst_line in function target_function_name? If so, provide a trace.
Output:
Query for Dependency Source Enumeration (Control Dependency) Below is your target code snippet.
<target code with line number>
Question: Which lines have control dependence over dst_line in function target_function_name? List all such lines.
Query for Pairwise Dependency Query (Information Flow) Below is your target code snippet.
<target code with line number>
Question: Is there information flow from (src_var, src_line) to (dst_var, dst_line) in function target_function_name? If so, provide a trace.
Output:
Query for Dependency Source Enumeration (Information Flow) Below is your target code snippet.
<target code with line number>
Question: Which variable instances have information flow over (dst_var, dst_line) in function target_function_name? List all such variables.
Output:
Appendix D Semantics-Aware Diverse Sampling
For each of the 180 annotated programs, we generate task instances for all three task types. For each type, up to five target variables (or lines) are selected per program. For each target, we sample up to five positive (if the specified dependency relationship holds) and five negative examples (otherwise) with enforced diversity, resulting in 12,553 task instances in total.
To construct our benchmark, we adopt a Semantics-Aware Diverse Sampling strategy that governs both target selection and task instance generation. The following sections describe how this sampling strategy is applied to select targets (Section D.1) and generate positive/negative task instances (Section D.2) for each task type.
D.1 Target Sampling
1
2Function SampleDataDepTargets($F$, $n$, $v_{t}$):
 Input: Function $F$, sample size $n$, initial target variable $v_{t}$
 Output: Set $T$ of sampled target variables
3
 $T\leftarrow\{v_{t}\}$ ;
 // Start with annotated target
 $S\leftarrow\mathtt{GetAllDataDepSources}(v_{t})$ ;
 // Track seen sources to reduce duplication
4
5  Shuffle all variables $V$ in $F$;
6
7 foreach $v\in V$ do
8    if $|T|\geq n$ then
9       break
10      end if
11
[…]
13       continue
14      end if
15
16    $T\leftarrow T\cup\{v\}$;
17    $S\leftarrow S\cup\mathtt{GetAllDataDepSources}(v)$;
18
19   end foreach
20 return $T$
21
22
230.5em Function IsValidDataDepTarget($v$):
24   stack $\leftarrow$ $(v,0,\emptyset)$];
25
26 while stack not empty do
27    $(curr,dist,visited\_sources)\leftarrow$ stack.pop();
28
29    foreach $src\in curr.data\_dep$ do
30       if $src\in visited\_sources$ then
          continue ;
          // Skip already visited
31
32         end if
33
34       $d\leftarrow dist+1$;
35       $S\leftarrow visited\_sources\cup\{src.name\}$;
36
37       if $d\geq 2$ or $|S|\geq 1$ then
38          return True
39         end if
40
41        Push $(src,d,S)$ to stack;
42
43      end foreach
44
45   end while
46 return False
47
Algorithm 1 Sample Data Dependency Task Instances
To construct a diverse and challenging benchmark, we sample up to five target variables (or lines) per program for each task type. A target serves as the anchor for generating task instances: positive and negative queries related to its dependencies. The sampling strategies are designed to enforce both diversity, by avoiding overlapping dependency traces, and complexity, by preferring targets involved in deeper or more nuanced dependency structures. While data dependency and information flow targets are sampled at the variable level based on transitive relationships, control dependency targets are line-based. Below, we describe the sampling procedures specific to each task type.
D.1.1 Data Dependency Target Sampling
Algorithm 1 describes how we sample target variables for data dependency tasks. This algorithm is used only for selecting target variables; positive and negative pairs are sampled separately and will be introduced later. Our goal is to ensure both diversity and complexity in the selected targets.
In function SampleDataDepTargets, given a program and an annotated initial target variable $v_{t}$, we aim to select up to $n=5$ target variables. The set $T$ is initialized with $v_{t}$ (line 2), and we maintain a set $S$ containing all transitive data-dependency sources of $v_{t}$ (Line 3). This helps avoid overlapping or redundant sampling. The function GetAllDataDepSources returns all variables that are transitively or directly data-dependent sources of the input variable.
Then, for each variable $v$ in the program (after random shuffling), we iterate through candidates and skip any that: (a) have already been selected, (b) appear in the transitive source set $S$ (i.e., are reachable from already-selected variables), (c) are not associated with a value assignment (i.e., reset is False), or (d) fail the structural filtering criteria defined in IsValidDataDepTarget (Line 9). In particular, if a variable $v$ is already in $S$, we exclude it because its data dependency sources would form a strict subset of those already covered. Any dependency trace from these sources to $v$ would also significantly overlap with traces of previously selected variables, reducing overall diversity.
We call IsValidDataDepTarget (lines 16–32) to ensure that the variable exhibits a non-trivial dependency structure. The function checks whether there exists a data-dependency source variable $v_{s}$ such that a transitive data dependency trace from $v_{s}$ to the input variable $v$ satisfies either of the following conditions:
  1. (a)
    The trace contains at least two edges (i.e., has a dependency depth $\geq$ 2), or
  2. (b)
    It involves at least one unique upstream variable name (excluding $v.name$ itself).
The function uses a DFS-style traversal initialized with a stack. Each element of the stack is a tuple containing three elements:
  • The current variable
  • Its distance from the input variable. Here, distance refers to the number of direct data dependency edges between the current variable and v, i.e., the length of the current transitiva data dependency trace. This value is used to determine whether the variable lies on a sufficiently deep trace (e.g., two or more edges), which is one of the criteria for accepting a variable as valid.
  • A set of source variable names already visited along the current path. The per-path tracking helps avoid revisiting nodes and prevents cycles during the depth-first traversal.
In each iteration (lines 18–31), we pop one tuple, and iterate over the direct data dependency sources of the current variable (line 20). We skip already visited sources to avoid cycles (lines 21–23). If the chain meets the success condition (line 26), we return True; otherwise, we continue exploring.
This approach ensures that each selected variable induces a trace that is sufficiently distinct and semantically meaningful.
D.1.2 Control Dependency Target Sampling
1
2
3Function SampleControlDepTargets($F$, $n$, $v_{0}$, depthdepth):
 Input: Function $F$, sample size $n$, annotated target variable $v_{0}$, max depth depthdepth
 Output: Set $T$ of selected target lines
4
5 $T\leftarrow\{v_{0}.line\}$;
 $L\leftarrow\texttt{GetControlDepLines}(v_{0},depth)$ ;
 // Blocked control-dep lines
6
7  Shuffle all variables $V$ in $F$;
8
9 foreach $v\in V$ do
10    if $|T|\geq n$ then
11       break
12
[…]
14       continue
15
16    $C\leftarrow\texttt{GetControlDepLines}(v,depth)$;
17    if $L\cap C\neq\emptyset$ then
       continue ;
       // Avoid overlapping control paths
18
19
20    $T\leftarrow T\cup\{v.line\}$;
21    $L\leftarrow L\cup C$;
22
23 return $T$
24
250.5em Function IsValidControlDepTarget($v$):
26   stack $\leftarrow$ [$(v,0)$];
27   visited $\leftarrow\{v.\mathtt{line}\}$;
28
29 while stack not empty do
30    $(curr,dist)\leftarrow$ stack.pop();
31
32    foreach $p\in curr.\mathtt{control\_dep}$ do
33       if $p.\mathtt{line}=curr.\mathtt{line}$ or $p.\mathtt{line}\in$ visited then
34          continue
35
36       $d\leftarrow dist+1$;
37       if $d\geq 2$ then
38          return True
39
40        visited $\leftarrow$ visited $\cup\{p.\mathtt{line}\}$;
41         Push $(p,d)$ to stack;
42
43
44 return False
Algorithm 2 Sample Control Dependency Target Variables
The control dependency sampling strategy (Algorithm 2 SampleControlDepTargets) follows the same template as data dependency target sampling but introduces line-based filtering to enforce structural diversity across control paths.
The key difference is that instead of tracking dependency sources, we track the set of control-dependency lines visited so far. Specifically, for each candidate variable $v$, we use GetControlDepLines to extract the line numbers associated with its control dependencies, up to a specified depth depthdepth. This is necessary because, in real-world code, certain lines, such as early return statements or large enclosing while loops, often dominate the control flow of the entire function. Without filtering, many variables would end up sharing identical or heavily overlapping control dependency traces, effectively blocking most of the lines before the loop even starts. By restricting overlap only within the first few levels of control dependency, we avoid such global blocking and enable the selection of variables that differ in their local control context, thus improving structural diversity in the sampled targets. If any of these lines overlap with those already visited, the candidate is skipped to avoid redundancy in control context (lines 11–12). The depth is set to 2.
The function IsValidControlDepTarget (lines 16–29) ensures that the variable is transitively control-dependent on at least two other lines (i.e., a path of length $\geq 2$). This is checked via depth-first traversal over control-dependency edges. Like before, only variables with meaningful and non-trivial dependency structure are retained.
D.1.3 Information Flow Target Sampling
1
2Function SampleInfoFlowTargets($F$, $n$, $v_{t}$):
 Input: Function $F$, sample size $n$, annotated target variable $v_{t}$
 Output: Set $T$ of selected target variables
3
4 $T\leftarrow\emptyset$;
 $S\leftarrow\emptyset$ ;
 // Track seen source variables across candidates
5
6  Shuffle all variables $V$ in $F$;
7
8 foreach $v\in V$ do
9    if $|T|\geq n-1$ then
10       break
11
[…]
13       continue
14
15    $U\leftarrow\mathtt{GetAllInfoFlowSources}(v,d)$;
16    $T\leftarrow T\cup\{v\}$;
17    $S\leftarrow S\cup U$;
18
 $T\leftarrow T\cup\{v_{t}\}$ ;
 // Add target after sampling
19 return $T$
20
210.5em
22 Function IsValidInfoFlowTarget($v$):
23 if IsValidDataDepTarget(v)(v) = False then
24    return False
25
26 $D\leftarrow\mathtt{GetAllDataDepSources}(v)$;
27 $D\leftarrow D\cup\{v\}$;
28
29 foreach $x\in D$ do
30    if $x.\mathtt{control\_dep}\neq\emptyset$ then
31       return True
32
33 return False
Algorithm 3 Sample Information Flow Target Variables
The sampling algorithm for information flow targets (Algorithm 3) extends the structure used in data dependency sampling (Section D.1.1 but differs in several key ways.
First, we do not initialize the selected set with the annotated target variable $v_{t}$ (line 2). Doing so would cause the seen source set $S$ to immediately block all annotated variables (see annotation process in Section 3.2). Instead, we defer adding $v_{t}$ until after sampling (line 13).
Second, the validity check is stricter. The function IsValidInfoFlowTarget (lines 15–23) requires that the variable meet the same structural criteria as in the data dependency case (line 16), and that at least one of its data-dependency sources (including itself) is control-dependent on at least one variable. This ensures that selected targets participate in implicit information flows.
Finally, the dependency sources are gathered using GetAllInfoFlowSources (line 10), which aggregates all upstream control and data dependencies. Aside from these differences, the traversal, blocking, and shuffling logic mirror those used in data dependency target sampling.
D.2 Positive and Negative Task Instances Sampling
For each selected target (a variable or line, depending on the task type), we construct up to five positive and five negative task instances to evaluate the model’s ability to identify program dependencies. Each instance consists of a query: whether a specified dependency relationship holds between two program elements, together with a ground-truth answer (boolean) with a trace (if the relationship holds). While the overall sampling strategy is consistent across task types, the specifics vary based on the dependency semantics: data dependency and information flow tasks operate at the variable level, while control dependency operates at the line level and is restricted to syntactic condition constructs. We use prioritized sampling strategies to ensure negative examples are both structurally plausible and semantically incorrect (i.e., the relationship does not hold), thereby maximizing evaluation difficulty. Below, we detail the sampling procedure for each task type.
D.2.1 Sampling Task Instances for Data Dependency
1Function SampleNegativeDataDepInstances($v_{t}$, $V$, OtherVarsOtherVars):
 Input: Target variable $v_{t}$, set of all variables $V$, set of irrelevant variables OtherVarsOtherVars
 Output: Two sets of variables: primary and secondary candidates
2
3 $D\leftarrow\mathtt{GetAllDataDepSources}(v_{t})$;
 $P\leftarrow\emptyset$ ;
 // Primary candidates: before target
 $S\leftarrow\emptyset$ ;
 // Secondary candidates: loop-local after target
4
5 foreach $x\in V\cup\mathtt{OtherVars}$ do
6    if $x==v_{t}$ or $x\in D$ or $x.\mathtt{reset}=\text{False}$ then
7       continue
8
9    if $x.\mathtt{line}<v_{t}.\mathtt{line}$ then
10       $P\leftarrow P\cup\{x\}$;
11
12    else if $\mathtt{InSameLoopBlock}(v_{t}.\mathtt{line},x.\mathtt{line})$ then
13       $S\leftarrow S\cup\{x\}$;
14
15
16 return $P,S$
Algorithm 4 Sample Negative Task Instances for Data Dependency
Given a selected target variable $v_{t}$, we sample up to five positive and five negative task instances for data dependency. Each instance is a variable that can form a query of the form “Does variable $x$ have data dependency over $v_{t}$?”, along with a ground-truth label and, if applicable, a valid dependency trace.
Positive Sampling.
Positive examples are sampled directly from the transitive data dependency sources of $v$ using GetAllDataDepSources (Algorithm 1 line 3). These variables are, by definition, those on which $v_{t}$ is data-dependent (either directly or transitively).
Negative Sampling.
Negative examples are chosen among variables that do not have any transitive or direct data dependency over $v_{t}$, but are structurally close enough to be potentially confusing.
We use the strategy shown in Algorithm 4 to identify two prioritized pools of negative candidates:
  • Primary Candidates ($P$): variables that appear on control lines before $v_{t}$ and do not have data dependency over $v_{t}$. These are likely to be mistaken as influences from upstream data.
  • Secondary Candidates ($S$): variables that appear after $v_{t}$ but are in the same loop body (line 10). This accounts for reverse-order influence that often arises in loops, while still ensuring no true data dependency exists.
We sample up to five negative instances, prioritizing from $P$ and using $S$ only when the primary pool is insufficient. This prioritized sampling helps create hard negatives that challenge models to distinguish structural correlation from semantics.
D.2.2 Sampling Task Instances for Control Dependency
1
2Function SampleNegativeControlDepInstances($v_{t}$):
3
 Input: Target variable $v_{t}$
 Output: Two lists of candidate condition lines: primary and secondary
4
 $C\leftarrow\mathtt{GetConditionBlocks}()$ ;
 // All condition blocks of the function
5 $D\leftarrow\mathtt{GetControlDepLines}(v_{t})$;
6
 $P\leftarrow\emptyset$ ;
 // Primary: non-controlling lines before target
 $S\leftarrow\emptyset$ ;
 // Secondary: non-controlling lines after target in same loop
7
8 foreach $cond\_line\in C$ do
9    if $cond\_line\in D$ then
10       continue
11
12    if $cond\_line<v_{t}.\mathtt{line}$ then
13       $P\leftarrow P\cup\{cond\_line\}$;
14
15    else if $cond\_line>v_{t}.\mathtt{line}$ and $\mathtt{InSameLoopBlock}(v_{t}.\mathtt{line},cond\_line)$ then
16       $S\leftarrow S\cup\{cond\_line\}$;
17
18
19 return $P,S$
Algorithm 5 Sample Negative Task Instances for Control Dependency
For control dependency, each task instance consists of a program and a pair of line numbers, forming a question of the form: “Does line $x$ have control dependency over line yy?”. The model must answer this question along with a supporting trace if the answer is positive. Both positive and negative candidates are sampled from condition lines only, such as those introduced by if, while, or for statements.
Positive Sampling.
[…]
Negative Sampling.
Negative candidates are sampled using Algorithm 5. Unlike arbitrary lines, we restrict negative sampling to lines that are syntactic conditions in the source code, ensuring that each candidate is a plausible controller. We extract all such condition blocks within a program (represented as $C$ in line 1) using a multi-lingual parser  [^4] https://github.com/PurCL/LLMSCAN implemented with Tree-sitter  [^5] https://github.com/tree-sitter/tree-sitter.
We organize the candidates into two prioritized groups:
  • Primary Candidates: condition lines that appear before the target line and do not have control dependency over it.
  • Secondary Candidates: condition lines that appear after the target but are in the same enclosing loop. These can be misleading because constructs like break, return, or exceptions may terminate execution early, making them appear influential.
We prioritize sampling from primary candidates and fall back to secondary candidates as needed to maintain a total of five negative instances.
D.2.3 Sampling Task Instances for Information Flow
For information flow, we follow the same structure as in data dependency task instance generation (Section D.2.1). Given a target variable $v_{t}$, we sample up to five positive and five negative instances.
Positive Sampling.
[…]
Negative Sampling.
Negative instances are sampled using the same algorithm described in Algorithm 4, but with the data dependency source set replaced by the information flow source set. Specifically, we substitute line 2 of the algorithm with: $D\leftarrow\texttt{GetAllInfoFlowSources}(v_{t})$.
D.3 Sampling Statistics
Table 6: CoRe task instance distribution
Task Type C/C++ Java Python Total
Data Dependency Pos. 510 564 740 1,814
Neg. 953 989 858 2,800
Total 1,463 1,553 1,598 4,614
Control Dependency Pos. 527 574 592 1,693
Neg. 628 680 651 1,959
Total 1,155 1,254 1,243 3,652
Information Flow Pos. 742 750 799 2,291
Neg. 641 688 667 1,996
Total 1,383 1,438 1,466 4,287
Total Pos. 1,779 1,888 2,131 5,798
Neg. 2,222 2,357 2,176 6,755
Total 4,001 4,245 4,307 12,553
Table 7: CoRe Lite task instance distribution
Task Type C/C++ Java Python Total
Data Dependency Pos. 69 61 79 209
Neg. 131 145 105 381
Total 200 206 184 590
Control Dependency Pos. 72 81 86 239
Neg. 85 86 79 250
Total 157 167 165 489
Information Flow Pos. 92 94 105 291
Neg. 75 73 65 214
Total 168 167 170 505
Total Pos. 233 236 270 739
Neg. 292 304 249 845
Total 525 540 519 1,584
Tables 7 and 7 show the comprehensive breakdown of task instances generated in CoRe and CoRe Lite. They report the distribution of positive and negative samples across all studied languages. In total, CoRe contains 12,553 task instances, while CoRe Lite includes 1,584.
D.4 Ablation Study of the Sampling Strategy
Table 8: Ablation Study: Semantics-Aware Sampling vs. Random Sampling
Language Semantics-Aware Sampling Random Sampling
C 82.84 83.86
Java 82.63 85.78
Python 81.61 85.06
F1 for Dependency Classification
Language Semantics-Aware Sampling Random Sampling
C 31.27 37.14
Java 31.47 37.67
Python 29.29 31.25
Correct Trace Rate for Trace Prediction
To investigate the impact of the semantics-aware sampling strategy, we conducted an experiment where we removed all heuristics and randomly sampled 400 task instances for the information flow tasks. The performance of Claude 3.5 on this randomly sampled dataset, compared to our original settings, is shown in Table 8.
As observed, performance is consistently better with random sampling. This indicates that random sampling tends to yield easier tasks, as it does not enforce the same level of reasoning complexity as our semantics-aware approach. It is important to note that different sampling strategies result in completely different datasets, and thus, a direct comparison of model performance across these datasets does not necessarily reflect a model’s inherent ability.
Appendix E Model Invocation Details
We invoke all models through their respective APIs or platforms using official or authorized access. Below are the details for each model:
  • Claude 3.7 Sonnet [5], Claude 3.5 Sonnet [4], LLaMA3.1 (405B) [42], and DeepSeek R1 [44] are accessed via AWS Bedrock  [^6] https://aws.amazon.com/bedrock through model IDs:
      Claude 3.7 Sonnet: us.anthropic.claude-3-7-sonnet-20250219-v1:0. We enable thinking mode with a thinking token budget of 1,024.
      Claude 3.5 Sonnet: anthropic.claude-3-5-sonnet-20240620-v1:0
      Llama 3.1 405B: meta.llama3-1-405b-instruct-v1:0
      DeepSeek R1: us.deepseek.r1-v1:0
  • DeepSeek V3 [58] is accessed through the official DeepSeek platform  [^7] https://www.deepseek.com/ with model ID deepseek-chat.
  • Gemini 2.5 Pro [25] is invoked via the Google’s Gemini Developer API [^8] https://ai.google.dev/gemini-api/docs, using model ID gemini-2.5-pro-preview-05-06.
  • GPT 4o [67], GPT o4-mini [68], and GPT o3 [68] are accessed through the OpenAI  [^9] https://platform.openai.com/docs/models, with model IDs gpt-4o, o4-mini, and o3.
  • Qwen3 (235B) [83] is accessed via OpenRouter  [^10] https://openrouter.ai/qwen/qwen3-235b-a22b with ID qwen/qwen3-235b-a22b.
All models are queried with a decoding temperature of 0 to minimize randomness, except GPT o3 and GPT o4-mini, which do not support zero temperature and are run with the only setting they support: temperature as one. The output length for all models is capped at 2,048 tokens.
Each model is expected to return a response in the predefined JSON format described in the prompts from Section J. If the response is ill-formed or unparsable, we re-query the model using a fallback prompt requesting only the structured output. We retry up to three times, after which the instance is marked as invalid if parsing still fails.
Fallback Prompt Your previous response could not be parsed correctly. Please re-read the prompt and ensure your answer strictly follows the required JSON format enclosed with ```<your response here>```.Ensure that your JSON is valid and matches the specification. Try again:
Appendix F Evaluation Metrics
F.1 Dependency Classification Metrics
For dependency classification, we report standard metrics: precision, recall, and F1 score, defined as follows:
Precision $=\frac{\text{TP}}{\text{TP}+\text{FP}}$
Recall $=\frac{\text{TP}}{\text{TP}+\text{FN}}$
F1 Score $=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
We define:
  • True Positive (TP): the model predicts that a dependency exists, and the instance is labeled positive.
  • False Positive (FP): the model predicts a dependency, but the instance is labeled negative.
  • False Negative (FN): the model predicts no dependency, but the instance is labeled positive.
Each task instance has a binary ground truth label indicating whether the specified dependency relation exists between the given pair of program elements.
F.2 Trace Evaluation Metrics
For task instances where the model predicts a dependency exists, it is also required to return a trace: a transitive sequence of direct dependencies (or flows) between program elements. We evaluate the trace at the edge level, where each edge is defined as a consecutive ordered pair (src $\rightarrow$ dst) in the predicted trace.
We classify each predicted edge into one of three categories:
  • Valid Edge (VE): The edge corresponds to a true direct dependency in the program.
  • Gap Edge: The source and target nodes are connected by a valid transitive path, but there is no direct edge src $\rightarrow$ dstin the program. This indicates that one or more intermediate nodes were skipped by the model (e.g., src $\rightarrow$ x $\rightarrow$ dst exists, but x is omitted).
  • Invalid Edge (IE): There is no valid dependency path, neither direct nor transitive, between src and dst.
For each predicted trace, we compute the following trace-level metrics:
$\text{VE}_{\text{trace}}$ $=\frac{\#\text{Valid Edges}}{\#\text{Total Edges in trace}}$
$\text{IE}_{\text{trace}}$ $=\frac{\#\text{Invalid Edges}}{\#\text{Total Edges in trace}}$
$\text{Missing Steps}_{\text{trace}}$ $=\sum_{e\in\text{gap edges}}\text{\# intermediate nodes skipped for e}$
The final reported metrics are the averages over all predicted traces:
Valid Edge Rate (VE) $=\frac{1}{N}\sum_{i=1}^{N}\text{VE}_{\text{trace}_{i}}$
Invalid Edge Rate (IE) $=\frac{1}{N}\sum_{i=1}^{N}\text{IE}_{\text{trace}_{i}}$
Average Missing Steps (Miss.) $=\frac{1}{N}\sum_{i=1}^{N}\text{Missing Steps}_{\text{trace}_{i}}$
Correct Trace Rate (CT) $=\frac{\#\text{Correct Traces}}{N}$
where $N$ is the total number of predicted traces, i.e., the number of task instances where the model predicts the dependency exists.
A perfect trace would consist only of valid direct edges, yielding $\text{VE}=100\%$, $\text{IE}=0\%$, and $\text{Miss.}=0$, and contributes to the numerator of CT.
F.3 Enumeration Task Metrics
In the enumeration task, the model is asked to list all program elements (variables or lines) that have a specific dependency relation (data dependency, control dependency, or information flow) over a given target element. This is a set prediction task, and we evaluate the model’s output using the following metrics:
  • Exact Match (EM): The prediction is considered correct only if the set of predicted elements exactly matches the ground truth set. Formally,
    $\text{EM}=\begin{cases}1&\text{if }\hat{S}=S\\ 0&\text{otherwise}\end{cases}$
    where $S$ is the ground-truth set, and $\hat{S}$ is the predicted set.
  • Precision (P): The fraction of predicted elements that are correct:
    $\text{P}=\frac{|\hat{S}\cap S|}{|\hat{S}|}$
  • Recall (R): The fraction of ground-truth elements that are correctly predicted:
    $\text{R}=\frac{|\hat{S}\cap S|}{|S|}$
  • F1 Score: The harmonic mean of precision and recall:
    $\text{F1}=\frac{2\cdot\text{P}\cdot\text{R}}{\text{P}+\text{R}}$
These metrics are computed per instance and then averaged across all enumeration queries in the benchmark. Exact match provides a strict correctness measure, while precision, recall, and F1 allow partial credit and provide more granular insight into model behavior.
Appendix G Complete Evaluation Results
Table 9: Complete dependency classification results reported as Precision, Recall, and F1 score (%).
Model Lang. CoRe CoRe Lite
Data Dependency Control Dependency Information Flow Data Dependency Control Dependency Information Flow
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Claude 3.7 C/C++ 64.76 80.00 71.58 89.12 88.61 88.87 91.74 71.83 80.57 67.90 79.71 73.33 88.00 91.67 89.80 95.77 73.91 83.44
Java 68.11 83.69 75.10 86.72 89.90 88.28 88.71 75.47 81.56 67.44 95.08 78.91 90.12 90.12 90.12 90.12 77.66 83.43
Python 76.49 79.59 78.01 84.82 86.82 85.81 91.61 68.34 78.28 74.16 83.54 78.57 91.03 82.56 86.59 90.00 68.57 77.84
Average 69.79 81.09 74.90 86.89 88.44 87.65 90.69 71.88 80.14 69.83 86.11 76.94 89.72 88.12 88.84 91.96 73.38 81.57
Claude 3.5 C/C++ 53.90 90.78 67.64 69.98 88.05 77.98 82.40 83.29 82.84 56.64 92.75 70.33 70.33 88.89 78.53 84.78 84.78 84.78
Java 54.80 90.07 68.14 65.62 91.46 76.42 80.69 84.67 82.63 52.21 96.72 67.82 69.81 91.36 79.14 85.71 82.98 84.32
Python 64.99 88.78 75.04 67.06 86.99 75.74 86.21 77.47 81.61 62.07 91.14 73.85 72.00 83.72 77.42 90.00 77.14 83.08
Average 57.90 89.88 70.27 67.55 88.83 76.71 83.10 81.81 82.36 56.97 93.54 70.67 70.71 87.99 78.36 86.83 81.63 84.06
DeepSeek R1 C/C++ 84.92 80.59 82.70 94.94 88.99 91.87 96.71 75.20 84.61 87.30 79.71 83.33 95.71 93.06 94.37 96.20 82.61 88.89
Java 84.07 81.38 82.70 96.67 90.94 93.72 92.37 72.67 81.34 84.13 86.89 85.48 95.00 93.83 94.41 94.37 71.28 81.21
Python 86.43 76.62 81.23 95.14 85.98 90.33 94.15 72.47 81.90 83.78 78.48 81.05 95.89 81.40 88.05 96.05 69.52 80.66
Average 85.14 79.53 82.21 95.58 88.64 91.97 94.41 73.45 82.62 85.07 81.69 83.29 95.53 89.43 92.28 95.54 74.47 83.59
DeepSeek V3 C/C++ 60.79 81.76 69.73 74.78 79.89 77.25 84.59 72.51 78.08 65.43 76.81 70.67 72.62 84.72 78.21 85.37 76.09 80.46
Java 63.81 84.40 72.67 73.61 83.10 78.07 82.83 69.47 75.56 60.87 91.80 73.20 77.27 83.95 80.47 85.92 64.89 73.94
Python 74.06 82.57 78.08 70.74 77.20 73.83 85.42 68.21 75.85 67.42 75.95 71.43 81.71 77.91 79.76 90.79 65.71 76.24
Average 66.22 82.91 73.49 73.04 80.06 76.38 84.28 70.06 76.50 64.57 81.52 71.77 77.20 82.19 79.48 87.36 68.90 76.88
Gemini 2.5 Pro C/C++ 81.36 94.12 87.27 86.36 97.34 91.53 95.89 94.34 95.11 81.01 92.75 86.49 86.42 97.22 91.50 90.91 97.83 94.24
Java 86.07 92.02 88.95 91.64 97.39 94.43 94.50 96.13 95.31 83.10 96.72 89.39 89.89 98.77 94.12 97.80 94.68 96.22
Python 87.90 93.24 90.49 89.22 95.10 92.07 94.55 95.62 95.08 86.05 93.67 89.70 91.86 91.86 91.86 92.59 95.24 93.90
Average 85.11 93.13 88.90 89.07 96.61 92.68 94.98 95.36 95.17 83.39 94.38 88.53 89.39 95.95 92.49 93.77 95.92 94.79
GPT o3 C/C++ 92.87 91.96 92.41 94.06 90.13 92.05 97.81 90.43 93.98 94.03 91.30 92.65 91.89 94.44 93.15 95.45 91.30 93.33
Java 93.51 94.50 94.00 96.35 91.99 94.12 94.53 92.13 93.32 95.16 96.72 95.93 93.75 92.59 93.17 94.38 89.36 91.80
Python 94.18 91.89 93.02 96.65 87.67 91.94 95.24 90.24 92.67 91.14 91.14 91.14 97.30 83.72 90.00 93.07 89.52 91.26
Average 93.52 92.78 93.14 95.69 89.93 92.70 95.86 90.93 93.32 93.44 93.05 93.24 94.31 90.25 92.11 94.30 90.06 92.13
GPT o4-mini C/C++ 89.86 76.47 82.63 94.09 87.67 90.77 95.85 74.80 84.03 94.64 76.81 84.80 93.06 93.06 93.06 96.10 80.43 87.57
Java 87.92 74.82 80.84 97.90 89.20 93.35 93.98 74.93 83.38 92.31 78.69 84.96 97.37 91.36 94.27 94.94 79.79 86.71
Python 92.26 74.05 82.16 96.34 84.46 90.01 96.17 72.34 82.57 91.18 78.48 84.35 97.22 81.40 88.61 95.83 65.71 77.97
Average 90.01 75.11 81.88 96.11 87.11 91.38 95.33 74.02 83.33 92.71 77.99 84.70 95.88 88.61 91.98 95.62 75.31 84.08
GPT 4o C/C++ 53.21 89.41 66.72 72.08 77.42 74.66 71.02 87.87 78.55 55.65 92.75 69.57 69.77 83.33 75.95 70.34 90.22 79.05
Java 57.03 87.77 69.13 71.08 81.36 75.87 66.95 83.47 74.30 51.89 90.16 65.87 70.97 81.48 75.86 69.30 84.04 75.96
Python 67.56 85.00 75.28 74.07 73.82 73.94 72.05 80.35 75.98 63.89 87.34 73.80 82.50 76.74 79.52 77.55 72.38 74.88
Average 59.27 87.39 70.38 72.41 77.53 74.82 70.01 83.90 76.28 57.14 90.08 69.75 74.41 80.52 77.11 72.40 82.21 76.63
Llama 3.1 405B C/C++ 46.09 79.80 58.44 63.16 79.70 70.47 73.05 74.53 73.78 48.39 86.96 62.18 66.28 79.17 72.15 74.47 76.09 75.27
Java 49.05 82.45 61.51 59.21 80.66 68.29 68.54 77.87 72.91 45.24 93.44 60.96 64.42 82.72 72.43 69.39 72.34 70.83
Python 56.98 78.92 66.18 59.04 76.69 66.72 70.33 73.59 71.93 56.76 79.75 66.32 61.39 72.09 66.31 76.64 78.10 77.36
Average 50.71 80.39 62.04 60.47 79.02 68.49 70.64 75.33 72.87 50.13 86.72 63.15 64.03 77.99 70.30 73.50 75.51 74.49
Qwen3 235B C/C++ 83.76 76.86 80.16 93.51 81.97 87.36 97.79 53.64 69.28 84.06 84.06 84.06 96.77 83.33 89.55 96.43 58.70 72.97
Java 82.66 72.70 77.36 93.70 85.54 89.44 92.99 54.80 68.96 84.21 78.69 81.36 96.05 90.12 92.99 96.15 53.19 68.49
Python 86.10 73.65 79.39 91.90 76.69 83.61 95.97 56.70 71.28 84.93 78.48 81.58 90.67 79.07 84.47 96.43 51.43 67.08
Average 84.17 74.40 78.97 93.04 81.40 86.80 95.58 55.05 69.84 84.40 80.41 82.33 94.50 84.17 89.00 96.34 54.44 69.51
Table 10: Complete trace evaluation results: Correct Trace Rate (CT, %), Valid Edge Rate (VE, %), Invalid Edge Rate (IE, %), and Avg. Missing Steps (Miss). $\uparrow$ means higher is better; $\downarrow$ lower is better.
Model Lang. CoRe CoRe Lite
Data Dependency Control Dependency Information Flow Data Dependency Control Dependency Information Flow
[…]
Claude 3.7 C/C++ 66.47 71.33 6.82 0.05 58.06 65.43 2.86 0.39 38.81 53.72 8.34 0.32 72.46 74.46 3.67 0.04 62.50 69.26 2.27 0.35 41.30 57.78 6.67 0.27
Java 60.64 69.43 9.85 0.11 59.06 68.36 3.84 0.39 35.20 55.59 10.51 0.40 72.13 77.62 7.90 0.16 62.96 73.27 2.04 0.33 39.36 59.16 10.31 0.35
Python 57.03 67.17 7.35 0.20 48.99 58.66 3.09 0.59 37.80 53.57 7.18 0.34 63.29 73.29 8.16 0.15 54.65 62.46 3.34 0.36 39.05 54.38 6.86 0.29
Average 61.38 69.31 8.01 0.12 55.37 64.15 3.26 0.46 37.27 54.29 8.68 0.35 69.29 75.12 6.58 0.12 60.04 68.33 2.55 0.35 39.90 57.11 7.95 0.30
Claude 3.5 C/C++ 50.39 65.67 23.38 0.06 55.79 67.98 9.51 0.24 31.27 55.29 15.67 0.52 55.07 69.79 20.91 0.06 56.94 70.80 6.20 0.28 40.22 61.47 11.42 0.57
Java 44.86 62.10 24.14 0.12 53.83 70.18 11.99 0.27 31.47 58.03 16.00 0.60 60.66 78.93 15.27 0.10 59.26 72.47 9.36 0.28 35.11 59.24 14.03 0.79
Python 42.30 60.11 24.62 0.20 48.99 65.11 9.86 0.40 29.29 52.20 14.84 0.53 46.84 64.77 22.03 0.25 55.81 72.23 5.21 0.28 32.38 55.76 11.42 0.37
Average 45.85 62.63 24.05 0.13 52.87 67.76 10.45 0.30 30.68 55.17 15.50 0.55 54.19 71.16 19.40 0.14 57.34 71.83 6.92 0.28 35.90 58.82 12.29 0.57
DeepSeek R1 C/C++ 71.96 74.78 5.11 0.01 67.17 76.17 1.06 0.31 41.64 55.69 6.68 0.40 69.57 71.50 7.49 0.01 73.61 83.75 1.39 0.28 40.22 58.48 4.17 0.53
Java 68.09 73.70 3.92 0.10 68.82 78.39 1.06 0.35 39.07 55.29 5.46 0.45 75.41 78.80 3.99 0.13 71.60 81.28 0.62 0.37 40.43 55.63 5.20 0.39
Python 57.57 65.92 3.84 0.27 57.94 69.55 0.82 0.51 35.79 53.79 5.05 0.54 56.96 67.85 4.14 0.34 54.65 66.24 0.39 0.42 38.10 52.76 5.43 0.45
Average 65.87 71.47 4.29 0.13 64.64 74.70 0.98 0.39 38.83 54.92 5.73 0.46 67.31 72.72 5.21 0.16 66.62 77.09 0.80 0.36 39.58 55.62 4.93 0.46
DeepSeek V3 C/C++ 56.27 63.99 13.91 0.07 43.83 55.80 7.77 0.38 22.64 41.10 21.65 0.39 60.87 66.11 5.99 0.07 48.61 63.43 7.27 0.35 21.74 41.58 24.09 0.37
Java 52.13 63.00 14.73 0.19 37.80 52.72 9.05 0.53 22.40 42.56 16.46 0.52 63.93 73.14 13.96 0.10 41.98 56.58 4.32 0.57 24.47 41.44 14.68 0.53
Python 45.41 57.03 14.57 0.43 32.09 43.88 7.52 0.63 19.65 37.10 19.29 0.60 44.30 54.99 12.71 0.35 30.23 40.81 7.40 0.78 24.76 40.41 14.91 0.50
Average 51.27 61.34 14.40 0.23 37.91 50.80 8.11 0.51 21.56 40.25 19.13 0.51 56.37 64.75 10.89 0.18 40.27 53.61 6.33 0.56 23.66 41.14 17.89 0.47
Gemini 2.5 Pro C/C++ 92.16 92.92 0.90 0.00 93.74 95.79 0.22 0.03 69.14 84.33 1.31 0.31 92.75 92.75 0.00 0.00 95.83 96.94 0.00 0.01 68.48 86.11 0.63 0.39
Java 83.16 87.82 2.81 0.06 92.51 95.28 0.06 0.07 68.40 85.26 1.38 0.37 88.52 93.85 2.32 0.03 92.59 95.23 0.00 0.11 67.02 83.40 0.35 0.39
Python 86.08 91.14 1.14 0.05 89.53 92.70 0.51 0.06 67.83 85.87 1.46 0.35 89.87 92.83 0.32 0.03 88.37 90.31 0.39 0.02 70.48 86.38 1.59 0.32
Average 87.13 90.63 1.62 0.04 91.93 94.59 0.26 0.05 68.46 85.15 1.38 0.34 90.38 93.14 0.88 0.02 92.26 94.16 0.13 0.05 68.66 85.30 0.86 0.37
GPT o3 C/C++ 89.61 90.47 0.85 0.01 79.70 84.63 0.60 0.12 52.56 72.21 3.70 0.42 88.41 88.89 0.97 0.01 86.11 90.00 0.00 0.12 56.52 76.56 0.72 0.41
Java 83.16 88.98 1.89 0.11 80.84 86.23 0.00 0.17 51.07 72.91 4.42 0.57 91.80 94.13 0.00 0.07 79.01 85.91 0.00 0.19 51.06 71.01 5.59 0.51
Python 78.92 86.81 0.79 0.16 71.28 78.54 0.11 0.26 49.19 72.08 3.43 0.57 78.48 87.04 1.27 0.16 67.44 74.77 0.00 0.21 49.52 72.17 3.42 0.52
Average 83.90 88.75 1.18 0.10 77.27 83.13 0.24 0.19 50.94 72.40 3.85 0.52 86.23 90.02 0.75 0.08 77.52 83.56 0.00 0.17 52.37 73.25 3.24 0.48
GPT o4-mini C/C++ 73.73 74.40 0.75 0.02 67.55 75.44 0.23 0.29 40.97 55.47 1.70 0.41 75.36 75.36 1.45 0.00 69.44 79.17 0.00 0.32 42.39 60.29 2.36 0.43
Java 62.77 67.35 2.26 0.13 67.94 76.18 0.41 0.37 41.07 55.79 2.71 0.48 73.77 75.41 0.00 0.10 74.07 81.54 0.41 0.31 43.62 58.92 2.77 0.62
Python 59.59 66.71 1.40 0.20 56.08 66.31 0.00 0.49 41.93 57.17 1.57 0.44 62.03 72.16 2.15 0.18 56.98 66.05 0.00 0.43 40.95 51.57 2.38 0.36
Average 65.36 69.49 1.47 0.12 63.86 72.64 0.21 0.38 41.32 56.14 1.99 0.44 70.39 74.31 1.20 0.09 66.83 75.59 0.14 0.35 42.32 56.93 2.50 0.47
GPT 4o C/C++ 60.39 66.18 15.66 0.12 41.56 51.95 8.49 0.38 23.72 44.56 29.31 0.66 60.87 65.16 17.45 0.14 43.06 56.81 5.76 0.51 26.09 49.96 24.29 0.70
Java 55.67 64.73 13.39 0.23 40.94 53.33 7.62 0.53 23.47 46.36 24.09 0.64 63.93 67.08 12.57 0.20 45.68 54.65 5.43 0.53 23.40 49.58 21.74 0.70
Python 52.30 61.90 13.01 0.33 35.47 47.07 5.51 0.55 20.15 41.81 24.43 0.82 60.76 70.53 6.22 0.42 40.70 54.26 5.23 0.56 18.10 38.18 22.26 0.83
Average 56.12 64.27 14.02 0.23 39.32 50.78 7.21 0.49 22.45 44.24 25.94 0.71 61.85 67.59 12.08 0.25 43.15 55.24 5.47 0.53 22.53 45.91 22.76 0.74
Llama 3.1 405B C/C++ 37.45 45.42 28.39 0.10 31.50 42.60 19.03 0.39 13.75 27.63 36.93 0.41 34.78 43.69 37.84 0.14 30.56 41.47 20.11 0.31 13.04 24.65 39.47 0.45
Java 32.80 44.21 30.06 0.21 26.13 39.13 17.96 0.56 15.73 33.16 31.98 0.58 50.82 61.17 25.30 0.18 27.16 38.17 13.95 0.60 14.89 33.31 27.63 0.64
Python 27.30 37.82 31.66 0.31 25.17 37.15 17.81 0.62 15.02 28.65 32.48 0.67 32.91 41.43 29.60 0.41 25.58 33.99 13.33 0.64 20.00 35.11 28.82 0.77
Average 32.52 42.48 30.04 0.21 27.60 39.63 18.27 0.52 14.83 29.81 33.80 0.55 39.50 48.76 30.91 0.24 27.77 37.88 15.80 0.52 15.98 31.02 31.97 0.62
Qwen3 235B C/C++ 68.43 70.74 4.85 0.02 59.39 68.04 2.12 0.30 29.78 39.82 4.78 0.27 73.91 76.81 7.25 0.00 61.11 73.10 1.62 0.28 31.52 44.58 2.97 0.32
Java 60.11 64.80 4.60 0.09 58.01 68.05 3.23 0.38 28.13 39.81 5.71 0.35 65.57 70.36 4.23 0.11 62.96 73.11 2.26 0.40 26.60 39.40 5.04 0.38
Python 55.27 63.43 4.03 0.24 48.14 58.02 1.62 0.49 28.66 41.59 5.96 0.35 56.96 65.62 6.75 0.33 53.49 60.66 0.68 0.48 31.43 38.91 6.65 0.23
Average 61.27 66.32 4.49 0.11 55.18 64.70 2.32 0.39 28.86 40.41 5.48 0.32 65.48 70.93 6.08 0.15 59.19 68.96 1.52 0.38 29.85 40.96 4.89 0.31
Table 11: Complete dependency source enumeration results reported as Exact Match Rate (EM), Precision (P), Recall (R), and F1 score (%).
Model Lang. CoRe Lite
Data Dependency Control Dependency Information Flow
EM P R F1 EM P R F1 EM P R F1
Claude 3.7 C/C++ 23.00 75.35 64.10 59.94 50.96 98.04 81.32 85.40 10.12 90.27 57.25 64.11
Java 16.99 75.07 60.90 57.31 59.28 97.62 87.01 90.21 3.59 91.93 59.56 67.39
Python 21.74 81.93 65.59 67.55 44.85 96.29 78.74 83.81 7.06 89.12 59.29 66.88
Average 20.58 77.45 63.53 61.60 51.70 97.32 82.36 86.47 6.92 90.44 58.70 66.13
Claude 3.5 C/C++ 13.50 55.01 73.95 57.89 40.13 89.65 79.49 80.81 10.71 84.37 61.15 67.12
Java 10.68 62.88 75.41 62.14 47.31 91.00 86.13 86.25 4.79 89.16 59.14 67.74
Python 9.24 64.06 60.47 56.39 33.94 90.08 80.21 81.68 7.65 87.40 58.70 66.15
Average 11.14 60.65 69.94 58.81 40.46 90.24 81.94 82.91 7.72 86.98 59.66 67.00
DeepSeek R1 C/C++ 46.50 90.24 75.35 77.84 56.05 98.68 80.59 85.84 8.93 92.03 48.93 59.63
Java 45.15 92.32 72.38 76.20 52.10 98.30 79.89 86.10 4.79 90.23 44.47 55.26
Python 25.00 89.35 59.01 65.67 36.97 97.80 72.57 80.27 7.65 93.41 41.64 52.70
Average 38.88 90.64 68.91 73.24 48.37 98.26 77.68 84.07 7.12 91.89 45.01 55.86
DeepSeek V3 C/C++ 20.00 64.24 66.86 58.53 26.11 79.61 68.37 69.31 3.57 66.60 62.59 60.09
Java 16.99 68.09 56.93 54.21 24.55 81.73 75.79 74.86 2.40 72.52 58.55 60.19
Python 10.87 71.06 56.89 57.29 13.33 76.35 62.97 65.21 1.76 75.08 56.90 59.69
Average 15.95 67.80 60.23 56.68 21.33 79.23 69.04 69.79 2.58 71.40 59.35 59.99
Gemini 2.5 Pro C/C++ 53.50 82.87 96.31 85.46 76.43 95.09 94.18 92.61 35.71 93.90 87.88 88.07
Java 43.69 82.89 95.41 84.80 79.64 96.08 96.22 95.12 20.36 92.84 88.55 88.45
Python 51.09 87.12 94.14 87.41 70.91 92.45 95.65 92.53 24.12 93.70 92.70 92.00
Average 49.43 84.29 95.29 85.89 75.66 94.54 95.35 93.42 26.73 93.48 89.71 89.51
GPT o3 C/C++ 53.00 89.54 91.63 88.55 68.79 98.30 87.41 90.16 18.45 95.14 78.84 84.39
Java 42.23 87.57 91.49 86.95 77.25 98.88 90.93 93.28 8.38 91.52 77.24 80.32
Python 30.43 88.64 90.08 87.05 66.67 98.34 87.80 90.94 20.00 96.62 79.50 84.43
Average 41.89 88.58 91.07 87.52 70.90 98.51 88.71 91.46 15.61 94.43 78.53 83.05
GPT o4-mini C/C++ 34.50 86.97 81.04 79.11 63.06 98.95 85.01 89.20 11.31 95.15 56.05 66.03
Java 35.44 89.01 71.52 71.65 66.47 99.64 86.46 90.67 5.99 95.05 50.31 61.39
Python 17.39 82.74 66.79 68.59 55.76 98.96 80.64 86.08 9.41 95.15 49.52 60.35
Average 29.11 86.24 73.12 73.12 61.76 99.18 84.04 88.65 8.90 95.12 51.96 62.59
GPT 4o C/C++ 14.00 64.52 51.70 48.65 31.85 86.53 70.07 74.10 4.17 73.60 46.18 53.19
Java 13.59 71.22 52.91 51.59 34.13 85.21 74.42 76.43 1.80 80.76 37.35 46.61
Python 9.24 74.97 48.58 53.95 21.21 85.17 70.74 73.23 2.94 80.85 35.18 44.90
Average 12.28 70.24 51.06 51.40 29.06 85.64 71.74 74.59 2.97 78.40 39.57 48.23
Llama 3.1 405B C/C++ 2.50 60.28 29.48 23.84 7.64 61.33 45.11 43.59 3.57 66.61 28.78 33.91
Java 5.83 59.15 29.52 26.21 5.99 64.25 42.85 42.25 1.20 67.67 25.31 30.83
Python 0.54 56.56 15.25 15.33 2.42 57.58 36.12 36.88 0.00 59.55 23.13 28.41
Average 2.96 58.66 24.75 21.79 5.35 61.05 41.36 40.91 1.59 64.61 25.74 31.05
Qwen3 235B C/C++ 25.50 95.93 34.94 34.83 47.77 96.84 70.24 73.82 5.95 97.87 21.23 25.28
Java 22.82 94.40 35.94 36.93 51.50 96.59 75.02 78.82 1.80 96.38 20.36 25.60
Python 15.76 95.58 28.14 28.97 34.55 95.13 63.34 66.90 6.47 96.96 18.55 22.88
Average 21.36 95.30 33.01 33.58 44.61 96.19 69.53 73.18 4.74 97.07 20.05 24.59
Complete evaluation results for CoRe and CoRe Lite are provided in Table 9 (dependency classification), Table 10 (trace generation), and Table 11 (dependency source enumeration).
CoRe vs. CoRe Lite
We observe minimal performance differences between CoRe and CoRe Lite, suggesting that CoRe Lite serves as a reliable alternative when computational resources are limited.
Performance Analysis on Task Types.
For trace generation 10, we observe that information flow tasks consistently exhibit significantly higher average missing steps (Miss.) and invalid edge (IE) rates compared to data and control dependency task types. This suggests that information flow is the most challenging task for LLMs, where they frequently generate incorrect edges. A comparison between Claude 3.5 and Claude 3.7 reveals that while their valid edge rates (VE) are similar, Claude 3.7 demonstrates a much lower invalid edge rate (IE). This indicates Claude 3.7 possesses better reasoning capabilities, leading to fewer erroneous edges.
Regarding dependency source enumeration (Table 11), different models display distinct error tendencies. For models such as Claude 3.7, DeepSeek R1, Llama 3.1 (405B), Qwen 3 (235B), GPT 4-mini, and GPT 4o, recall is consistently lower than precision across all task types (data dependency, control dependency, and information flow). This pattern suggests these models tend to yield precise dependency sources but frequently miss many relevant ones. Conversely, models like Gemini 2.5 Pro and DeepSeek V3 do not exhibit such a consistent trend in their precision-recall balance.
Performance Analysis on Programming Languages.
Overall, we do not observe a significant performance trend across programming languages for dependency classification (Table 9) or trace generation (Table 10). This holds across task types (data dependency, control dependency, and information flow) and models, with no evidence that any language is consistently more challenging or that any model is particularly strong in one language. This is likely because performance is heavily influenced by the specific structure of each program.
However, for the dependency source enumeration task (Table 11), we observe a consistent pattern in most of the cases: models perform notably better on C and Java than on Python, with performance gaps sometimes substantial. We attribute this to the higher information density of Python due to the lack of explicit block delimiters (e.g., {}) and the use of syntactic sugar. The absence of typing annotations makes precise reasoning more difficult.
Appendix H The Impact of Different Experimental Settings on Model Behaviors
(a) F1 scores for dependency classification
(b) Correct trace rates for trace generation
Figure 4: Impact of query type (classification only vs. with trace) and input clipping on F1 score and correct trace rate across tasks.
(a) F1 scores for dependency classification
(b) Correct trace rates for trace generation
Figure 5: F1 and correct trace rate under FSL as the number of retrieved examples ($k$) increases.
We conduct a few-shot learning (FSL) study using both sparse (BM25) and dense (CodeBERT [34]) retrievers to retrieve the top-$k$ most similar examples from the rest of the entire dataset, with $k\in\{4,8,12,16,20\}$. Retrieved examples are sorted by similarity score, with the most relevant placed closest to the query in the prompt. We evaluate this setup using Claude 3.7 in thinking mode, focusing on the pairwise dependency classification task. For each query, the model is asked whether the dependency holds, and if so, to generate a valid trace. To mitigate label imbalance and reduce bias, we ensure an equal number of positive and negative examples in each prompt.
As shown in Fig. 5, performance drops as the number of retrieved few-shot examples ($K$) increases for both dependency classification and trace generation. This degradation supports our hypothesis that longer inputs dilute model attention from task instructions and that retrieved examples, despite being label-balanced, can introduce bias. Furthermore, standard retrieval methods (e.g., BM25, CodeBERT) struggle to capture task-relevant structural and semantic similarity, highlighting the need for more retrieval strategies tailored to code reasoning.
Appendix I Other experimental explorations
Table 12: Effect of adding reverse dependency reasoning examples to the prompts on dependency classification and trace generation tasks.
Language Original Modified Prompt
C 67.64 66.86
Java 68.14 68.34
Python 75.04 74.60
F1 for Dependency Classification
Language Original Modified Prompt
C 50.39 50.00
Java 44.86 45.57
Python 42.30 39.32
Correct Trace Rate for Trace Generation
I.1 Prompt with Reverse Dependency Examples
In Section 5.2, the experimental results shows that LLMs struggle with queries that involve reverse dependency order. To further explore the capabilities of LLMs in dealing with such complex dependencies, we added an example of reverse dependency reasoning to the prompt and conducted experiments with Claude 3.5 on data dependency tasks. The results are shown in Table 12.
The results show negligible differences with no significant improvement or decrease. However, our primary contribution is providing a robust dataset rather than optimizing model performance through prompt engineering. Including examples or things to pay attention to for all challenging scenarios would be impractical and could be considered a form of overfitting to specific cases rather than evaluating genuine reasoning capabilities.
I.2 Shorten the Prompt
Table 13: Effect of prompt length reduction on dependency classification and trace generation tasks.
Language Original Shortened
C 67.64 67.50
Java 68.14 68.17
Python 75.04 74.04
F1 for Dependency Classification
Language Original Shortened
C 50.39 49.80
Java 44.86 46.63
Python 42.30 39.86
Correct Trace Rate for Trace Generation
Our prompts (section J) are designed to be concise while including necessary definitions and illustrative examples, which are particularly important for non-reasoning models and align with common few-shot learning practices [15]. Modern LLMs also support very long input contexts, making prompt length less of a constraint.
To investigate the impact of prompt length, we conducted an experiment where we shortened the prompts by cutting half of the examples. The results for Claude 3.5 on data dependency tasks are in Table 13. These results show negligible overall differences, with a slight decrease in some metrics.
Appendix J Prompt Design
In this section, we include the prompt templates used for both pairwise dependency queries and target-centric dependency source enumeration (Section 3.3.1, covering data dependency, control dependency, and information flow. For clarity, content dynamically generated per data point, such as function bodies or variable names, is shown in blue, while the rest represents the fixed prompt structure.
Prompt for Pairwise Dependency Query (Data Dependency) You are a program analysis assistant. Perform a static data dependence analysis on a given code snippet, treating each branch or loop condition as potentially taking any outcome, without using semantic or symbolic execution to prune paths. ## 1. Data Dependence Definition Data dependence captures the influence of data flow between variables. We denote each variable instance as (var,lineNumber), meaning the variable var defined or updated at lineNumber. Direct Data Dependence
A variable instance (varB,lineB) is directly data dependent on (varA,lineA) if lineA writes to varA. Then, without overwritting the value of varA, lineB reads varA and computes the value of varB. In other words, the value of (varB,lineB) relies on the value produced at (varA,lineA). In addition, changing the execution order of (varA,lineA) and (varB,lineB) would alter their results.
A variable instance (varB,lineB) is data dependent on (varA,lineA) if there is a transitive (indirect) chain of direct data dependencies from (varA,lineA) to (varB,lineB). This is equivalent to saying that (varA,lineA) has data dependence over (varB,lineB), meaning the value written at (varA,lineA) can propagate to (varB,lineB) through a sequence of read/write operations. ## 2. Output Format When asked “Does (varA,lineA) have data dependence over (varB,lineB)? If so, provide one feasible trace.“: If data dependence exists:
‘‘‘json
"DataDependence": true,
"Trace": [
{ "from": ["varA", lineA], "to": ["varX", lineX] },
{ "from": ["varY", lineY], "to": ["varB", lineB] }
If no data dependence exists:
‘‘‘json
"DataDependence": false
A trace represents a transitive data flow, where each item is a direct data dependence edge. The full trace should begin with (varA,lineA) and end with (varB,lineB). Only one valid chain is required even if multiple chains are possible. ## 3. Intraprocedural Data Dependence All dependence analysis is performed within a single function. We do not track dependencies across function boundaries. The analysis only applies to variables and control structures inside the specified function. ## 4. Example Code Snippet ### Example 1 ‘‘‘python
1 total = 0
2 value = 1
3 step = value
4 if step > 1:
5 value += 3
6 while total <= 10:
7 total += value
8 difference = total - step
9 step += 1
10 final_result = total * 2
#### Example Question 1.1:
Does (value,2) have data dependence over (step,9)? If so, provide a trace.
Analysis:
– Line 9: step += 1 $\Rightarrow$ (step,9) directly depends on (step,3) (or itself).
– Line 3: step = value $\Rightarrow$ (step,3) directly depends on (value,2).
Hence the chain (value,2) $\rightarrow$ (step,3) $\rightarrow$ (step,9).
Output:
‘‘‘json
"DataDependence": true,
"Trace": [
{ "from": ["value", 2], "to": ["step", 3] },
{ "from": ["step", 3], "to": ["step", 9] }
#### Example Question 1.2:
Does (value,2) have data dependence over (total,7)? If so, provide a trace.
Analysis:
– Line 7: total += value reads value from (value,2) or (value,5).
– Line 5: value += 3 depends on (value,2) but may be conditionally executed.
Providing one feasible chain: (value,2) $\rightarrow$ (value,5) $\rightarrow$ (total,7).
Output:
‘‘‘json
"DataDependence": true,
"Trace": [
{ "from": ["value", 2], "to": ["value", 5] },
{ "from": ["value", 5], "to": ["total", 7] }
#### Example Question 1.3:
Does (step,3) have data dependence over (final_result,10)? If so, provide a trace.
Analysis: No direct or transitive dependence exists. Output:
‘‘‘json
"DataDependence": false
### Example 2 ‘‘‘python
1 arr = [1, 2, 3]
2 x = 5
3 if x > 2:
4 arr[0] = x
5 for i in range(3):
6 arr[i] += 1
7 temp = arr[i]
8 y = temp + 1
9 result = arr[-1]
#### Example Question 2.1:
Does (x,2) have data dependence over (arr,6)? If so, provide a trace.
Analysis: One feasible chain is (x,2) $\rightarrow$ (arr,4) $\rightarrow$ (arr,6). Output:
‘‘‘json
"DataDependence": true,
"Trace": [
{ "from": ["x", 2], "to": ["arr", 4] },
{ "from": ["arr", 4], "to": ["arr", 6] }
#### Example Question 2.2:
Does (i,5) have data dependence over (temp,7)? If so, provide a trace.
Analysis: Direct dependence (i,5) $\rightarrow$ (temp,7). Output:
‘‘‘json
"DataDependence": true,
"Trace": [
{ "from": ["i", 5], "to": ["temp", 7] }
#### Example Question 2.3:
Does (i,5) have data dependence over (result,9)? If so, provide a trace.
Analysis: Feasible chain (i,5) $\rightarrow$ (arr,6) $\rightarrow$ (result,9). Output:
‘‘‘json
"DataDependence": true,
"Trace": [
{ "from": ["i", 5], "to": ["arr", 6] },
{ "from": ["arr", 6], "to": ["result", 9] }
[YOUR TURN] Below is your target snippet. “‘
<target code with line number>
Question: Does (src_var, src_line) have data dependence over (dst_var, dst_line) in function target_function_name? If so, provide a trace.
Output:
Prompt for Dependency Source Enumeration (Data Dependency) You are a program analysis assistant. Perform a static data dependence analysis on a given code snippet, treating each branch or loop condition as potentially taking any outcome, without using semantic or symbolic execution to prune paths. ## 1. Data Dependence Definition Data dependence captures the influence of data flow between variables. We denote each variable instance as (var,lineNumber), meaning the variable var defined or updated at lineNumber. Direct Data Dependence
A variable instance (varB,lineB) is directly data dependent on (varA, lineA) if lineA writes to varA. Then, without overwritting the value of varA, lineB reads varA and computes the value of varB. In other words, the value of (varB,lineB) relies on the value produced at (varA,lineA). In addition, changing the execution order of (varA,lineA) and (varB,lineB) would alter their results.
A variable instance (varB,lineB) is data dependent on (varA,lineA) if there is a transitive (indirect) chain of direct data dependencies from (varA,lineA) to (varB,lineB). This is equivalent to saying that (varA,lineA) has data dependence over (varB,lineB), meaning the value written at (varA,lineA) can propagate to (varB,lineB) through a sequence of read/write operations. ## 2. Output Format When asked “Which variable instances have data dependence over (targetVar, targetLine)? List all such variables.”, respond:
‘‘‘json
"DataDependenceSources": [
["valA", lineA],
["valB", lineB],
If you believe there are no variables with data dependence over (targetVar, targetLine), respond:
‘‘‘json
"DataDependenceSources": [ ]
## 3. Intraprocedural Data Dependence All dependence analysis is performed within a single function. We do not track dependencies across function boundaries. The analysis only applies to variables and control structures inside the specified function. ## 4. Example Code Snippet ### Example 1 ‘‘‘python
1 total = 0
2 value = 1
3 step = value
4 if step > 1:
5 value += 3
6 while total <= 10:
7 total += value
8 difference = total - step
9 step += 1
10 final_result = total * 2
#### Example Question 1.1:
Which variable instances have data dependence over (step,9)? List all such variables.
Analysis:
– Line 9: step += 1 $\Rightarrow$ (step,9) directly depends on (step,3) (or on itself).
– Line 3: step = value $\Rightarrow$ (step,3) directly depends on (value,2).
Hence the transitive chain is (value,2) $\rightarrow$ (step,3) $\rightarrow$ (step,9).
Output:
‘‘‘json
"DataDependenceSources": [
["step", 9],
["step", 3],
["value", 2],
#### Example Question 1.2:
Which variable instances have data dependence over (total,7)? List all such variables.
Analysis:
– Line 7: total += value reads total (from line 1 or itself) and value (from line 2 or 5).
– Line 5: value += 3 depends on (value,2) but is under an if guard; static analysis assumes both execution paths.
Therefore possible sources are (total,1), (total,7), (value,5), and (value,2).
Output:
‘‘‘json
"DataDependenceSources": [
["total", 1],
["total", 7],
["value", 5],
["value", 2]
### Example 2 ‘‘‘python
1 arr = [1, 2, 3]
2 x = 5
3 if x > 2:
4 arr[0] = x
5 for i in range(3):
6 arr[i] += 1
7 temp = arr[i]
8 y = temp + 1
9 result = arr[-1]
#### Example Question 2.1:
Which variable instances have data dependence over (arr,6)? List all such variables.
Analysis:
– Line 6 modifies arr[i]. If i == 0, it relies on (arr,4); otherwise on (arr,1).
– Line 4 depends on (x,2).
i itself originates at line 5.
Output:
‘‘‘json
"DataDependenceSources": [
["arr", 6],
["arr", 4],
["arr", 1],
["i", 5],
["x", 2]
#### Example Question 2.2:
Which variable instances have data dependence over (temp,7)? List all such variables.
Analysis:
– Line 7: temp = arr[i] $\Rightarrow$ depends on arr[i] (from line 6) and on i (line 5).
Output:
‘‘‘json
"DataDependenceSources": [
["arr", 6],
["i", 5],
#### Example Question 2.3:
Which variable instances have data dependence over (result,9)? List all such variables.
Analysis:
– Line 9: result = arr[-1] where arr[-1] (arr[2]) may come from initialization (line 1) or updates (line 6), which in turn depend on i (line 5).
Output:
‘‘‘json
"DataDependenceSources": [
["arr", 6],
["arr", 1],
["i", 5],
[YOUR TURN] Below is your target snippet. “‘
<target code with line number>
Question: Which variable instances have data dependence over (dst_var, dst_line) in function target_function_name? List all such variables.
Output:
Prompt for Pairwise Dependency Query (Control Dependency) You are a program-analysis assistant. Your task is to statically analyze the control dependence of a given code snippet. ## 1. Control Dependence Definition Control dependence captures the influence of control flow decisions on the execution of statements. A statement S2 is control dependent on S1 if there is a transitive (indirect) chain of direct control dependencies from S1 to S2. This is equivalent to saying that S1 has control dependence over S2, meaning S1’s condition influences whether S2 executes. Direct Control Dependence:
A statement S2 is directly control dependent on a statement S1 if
1.  S1 is a conditional control statement (e.g. an if, while, for, switch, etc.).
2.  S1 directly determines whether S2 executes. That is, S1 has multiple successor branches,
there exists at least one branch in which S2 always executes, and there exists at least one other branch in which S2 does not necessarily execute. “Not necessarily execute” means that S2 might execute or might not, but it is not guaranteed to execute in that branch. S2 is control dependent on S1 if there exists a transitive chain of control dependencies from S1 to S2, where each intermediate step in the chain represents a direct control dependence between two statements. ## 2. Output Format When asked, “Does line S1 have control dependence over line S2? If so, provide a trace.” you should respond in JSON format as follows: If there is a control dependence:
‘‘‘json
"ControlDependence": true,
"Trace": [S1, ..., S2]
If there is no control dependence:
‘‘‘json
"ControlDependence": false
A trace represents a transitive control flow, where each adjacent pair in the list reflects a direct control dependence. The trace must start with S1 and end with S2. ## 3. Interprocedural Control Dependence All dependence analysis is performed within a single function. We do not track dependencies across function boundaries. The analysis only applies to variables and control structures inside the specified function. ## 4. Example Code Snippet ### Example 1 ‘‘‘python
1 if x > 0:
2 y = 10 # this line is directly control-dependent on line 1 (x>0)
3 if y > 5:
4 z = 20 # this line is directly control-dependent on line 3 (y>5)
5 w = 30 # this line is directly control-dependent on line 3 (y>5)
6 v = 40 # this line is NOT control-dependent on any line
Analysis:
– Line 2 is directly control dependent on line 1 (x>0).
– Lines 4 and 5 are directly control dependent on line 3 (y>5).
– Lines 4 and 5 are
indirectly control dependent on line 1 (x>0) because even if line 1 (x>0) is true, lines 4 and 5 may not execute if line 3’s condition (y>5) is false. However, since control dependence is transitive, lines 4 and 5 are indirectly control dependent on line 1.
– Line 6 is not control dependent on any lines because it always executes, regardless of whether line 1 (x>0) or line 3 (y>5) is true or false.
#### Example Question 1.1:
Does line 1 have control dependence over line 5? If so, provide a trace.
Output:
‘‘‘json
"ControlDependence": true,
"Trace": [1, 3, 5]
#### Example Question 1.2:
Does line 3 have control dependence over line 6? If so, provide a trace.
Output:
‘‘‘json
"ControlDependence": false
### Example 2 ‘‘‘python
1 count = 0
2 if count < 5:
3 count += 1
4 print("Step 1 done")
5 while count < 10:
6 if count == 7:
7 break
8 if count % 2 == 0:
9 continue
10 count += 2
11 if count > 9:
12 count = 9
13 print("Iteration done")
14 print("End of program")
#### Example Question 2.1:
Does line 5 have control dependence over line 13? If so, provide a trace.
Analysis:
– Line 13 is directly control dependent on line 8 because if line 8 evaluates to true, the continue skips line 13 for that iteration.
– Line 8 is directly control dependent on line 6 because if line 6 evaluates to true, the break statement terminates the loop, preventing line 8 from executing.
– Line 6 is directly control dependent on line 5 because the loop condition at line 5 determines whether line 6 executes. If line 5 evaluates to false, execution skips the loop entirely.
[…]
Output:
‘‘‘json
"ControlDependence": true,
"Trace": [5, 6, 8, 13]
#### Example Question 2.2:
Does line 8 have control dependence over line 10? If so, provide a trace.
Analysis: Similar to the explanation above, line 8 is an if condition inside the while loop with a continue, so it can skip the rest of the loop entirely, including line 10. Therefore, line 10 is directly control dependent on 8. Output:
‘‘‘json
"ControlDependence": true,
"Trace": [8, 10]
#### Example Question 2.3:
Does line 5 have control dependence over line 14? If so, provide a trace.
Analysis: Line 5 does not affect whether line 14 is reached. Line 14 is outside the while loop (lines 5–13) and will be executed regardless of the condition. Output:
‘‘‘json
"ControlDependence": false
#### Example Question 2.4:
Does line 2 have control dependence over line 12? If so, provide a trace.
Analysis: Line 2 does not affect whether line 12 is reached. The condition in line 2 only controls whether line 3 is executed. Line 12 is inside the while loop (lines 5–13) and may or may not execute regardless of the condition in line 2. Output:
‘‘‘json
"ControlDependence": false
[YOUR TURN] Below is your target snippet. “‘
<target code with line number>
Question: Does src_line have control dependence over dst_line in function target_function_name? If so, provide a trace.
Output:
Prompt for Dependency Source Enumeration (Control Dependency) You are a program-analysis assistant. Your task is to statically analyze the control dependence of a given code snippet. ## 1. Control Dependence Definition Control dependence captures the influence of control-flow decisions on the execution of statements. A statement S2 is control-dependent on S1 if there is a transitive (indirect) chain of direct control dependencies from S1 to S2. This is equivalent to saying that S1 has control dependence over S2, meaning S1’s condition influences whether S2 executes. Direct Control Dependence: A statement S2 is directly control-dependent on a statement S1 if: 1. S1 is a conditional control statement (e.g., an if, while, for, switch, etc.). 2. S1 directly determines whether S2 executes. That is, S1 has multiple successor branches, - there exists at least one branch in which S2 always executes, and - there exists at least one other branch in which S2 does not necessarily execute. “Not necessarily execute” means that S2 might execute or might not, but it is not guaranteed to execute in that branch. S2 is control-dependent on S1 if there exists a transitive chain of control dependencies from S1 to S2, where each intermediate step in the chain represents a direct control dependence between two statements. ## 2. Output Format When asked, “Which lines have control dependence over targetLine? List all such lines.” you should respond in JSON format as follows:
‘‘‘json
"ControlDependenceSources": [S1, ..., S2]
If you believe there are no lines that have control dependence over targetLine, respond:
‘‘‘json
"ControlDependenceSources": []
## 3. Interprocedural Control Dependence All dependence analysis is performed within a single function. We do not track dependencies across function boundaries. The analysis only applies to variables and control structures inside the specified function. ## 4. Example Code Snippet ### Example 1 python
1 if x > 0:
2 y = 10 # this line is directly control-dependent on line 1 (x>0)
3 if y > 5:
4 z = 20 # this line is directly control-dependent on line 3 (y>5)
5 w = 30 # this line is directly control-dependent on line 3 (y>5)
6 v = 40 # this line is NOT control-dependent on any line
Analysis: - Line 2 is directly control-dependent on line 1 (x>0) - Lines 4 and 5 are directly control-dependent on line 3 (y>5). - Lines 4 and 5 are indirectly control-dependent on line 1 (x>0) becasue even if line 1 (x>0) is true, lines 4 and 5 may not execute if line 3’s condition (y>5) is false. However, since control dependence is transitive, lines 4 and 5 are indirectly control-dependent on line 1. - Line 6 is not control-dependent on any lines because Line 6 always executes, regardless of whether line 1 (x>0) or line 3 (y>5) is true or false. #### Example Question 1.1: Which lines have control dependence over line 5? List all such lines. Output:
‘‘‘json
"ControlDependenceSources": [1, 3]
#### Example Question 1.2: Which lines have control dependence over line 6? List all such lines. Output:
‘‘‘json
"ControlDependenceSources": []
### Example 2 python
1 count = 0
2 if count < 5:
3 count += 1
4 print("Step 1 done")
5 while count < 10:
6 if count == 7:
7 break
8 if count % 2 == 0:
9 continue
10 count += 2
11 if count > 9:
12 count = 9
13 print("Iteration done")
14 print("End of program")
#### Example Question 2.1: Which lines have control dependence over line 13? List all such lines. Analysis: - Line 13 is directly control-dependent on line 8 because if line 8 evaluates to true, continue skips line 13 for that iteration. - Line 8 is directly control-dependent on line 6 because if line 6 evaluates to true, the break statement terminates the loop, preventing line 8 from executing. - Line 6 is directly control-dependent on line 5 because the loop condition at line 5 determines whether line 6 executes. If line 5 evaluates to false, execution skips the loop entirely. Output:
‘‘‘json
"ControlDependenceSources": [5, 6, 8]
#### Example Question 2.2: Which lines have control dependence over line 10? List all such lines. Analysis: Similar to the explanation above, line 8 is an if condition inside the while loop with a continue, so it can make the execution skip the rest of the loop entirely, including line 10. Output:
‘‘‘json
"ControlDependenceSources": [5, 6, 8]
#### Example Question 2.3: Which lines have control dependence over line 14? List all such lines. Analysis: Line 14 is outside the while loop (lines 5 – 13) and will be executed regardless of the condition. Output:
‘‘‘json
"ControlDependenceSources": []
#### Example Question 2.4: Which lines have control dependence over line 12? List all such lines. Analysis: Line 2 doesn’t affect whether line 12 is reached. The condition in line 2 only controls whether line 3 is executed. Line 12 is in the while loop (lines 5 – 13) and will or will not be executed regardless of the condition in line 2. Output:
‘‘‘json
"ControlDependenceSources": [5, 6, 8, 11]
[YOUR TURN] Below is your target snippet. “‘
<target code with line number>
Question: Which lines have control dependence over dst_line in function target_function_name? List all such lines.
Output:
Prompt for Pairwise Dependency Query (Information Flow) You are a program-analysis assistant. Please perform a static analysis of “information flow” on a given code snippet, treating each branch or loop condition as potentially taking any outcome, without using semantic or symbolic execution to prune paths. ## 1. Information Flow Definition An information flow from variable x to variable y, written x -> y, exists whenever information stored in x is transferred to or used to derive information transferred to y. We denote each variable instance as (var,lineNumber), meaning the variable var is defined or updated at lineNumber. ### Types of Direct Flows Direct Explicit Flow
A direct explicit flow occurs when the value written by one variable instance is directly used in computing the value of another, through operations like assignments.
Formally, if (varA, lineA) writes a value and (varB, lineB) reads that value to compute its own, then there is a direct explicit flow from (varA, lineA) to (varB, lineB), written as: (varA,lineA) -> (varB,lineB). Example 1:
‘‘‘python
1 x = 5
2 y = x + 1
(x,1) -> (y,2) is a direct explicit flow. Example 2:
‘‘‘python
1 x = 5
2 if cond:
3 y = x + 1
(x,1) -> (y,3) is a direct explicit flow. Under static analysis, we treat each conditional or loop as potentially taking any outcome, and we do not perform semantic pruning. Therefore, even if the statement at line 3 is conditionally executed, we still recognize (x,1) as used to compute (y,3).
Direct Implicit Flow
A direct implicit flow from (varA,lineA) to (varB, lineB) occurs when (varA,lineA)’s value directly influences whether lineB executes. In other words, it corresponds to a direct control dependence, where the condition has at least one branch where (varB, lineB) must run and another where it may not. All conditional structures (e.g. if, while, for, switch, etc.) generate these flows.
Formally, if (varA,lineA) is directly read in a condition at lineC and that condition directly decides whether lineB executes, then we say there is a direct implicit flow from (varA,lineA) to (varB,lineB) (written (varA,lineA) -> (varB,lineB)). Example:
‘‘‘python
1 z = x + 5
2 if z > 1:
3 y = 10
(z,1) -> (z,2) -> (y,3) is a direct implicit flow through the condition at line 2. The asterisk  indicates that the variable is used for the conditional statement but not redefined at that line (i.e. used in a condition).
### Information Flow Between Variables An information flow exists from a variable (varA, lineA) to another variable (varB, lineB) if there is a transitive chain of direct flows — explicit, implicit, or both — connecting them: (varA, lineA) -> ... -> (varB, lineB) ## 2. Output Format When asked “Is there information flow from (varA,lineA) to (varB,lineB)? If so, provide one feasible trace.”, respond with a JSON object: If information flow exists:
‘‘‘json
"InformationFlow": true,
"Trace": [
"from": ["varOrExpr", lineNumber /*, "use" optional */],
"to": ["varOrExpr", lineNumber /*, "use" optional */],
"type": "data" | "control"
If no information flow exists:
‘‘‘json
"InformationFlow": false
Every trace edge must specify "from", "to" (each may include a third "use" element when the line is a conditional statement that only reads without redefining/updating the value), and "type". We do not consider loop in the trace. A trace represents a transitive chain of direct flows arising from compositions of direct implicit or explicit flows, where each edge represents a direct implicit flow (through control dependence) or a direct explicit flow (through data dependence). You only need to provide one valid trace if you conclude an information flow exists, even if multiple possible chains exist. ## 3. Intraprocedural Data Dependence All dependence analysis is performed within a single function. We do not track dependencies across function boundaries. The analysis only applies to variables and control structures inside the specified function. ## 4. Example Code Snippet ### Example 1 ‘‘‘python
1 def example_func():
2 status = 0
3 flag = False
4 balance = 1000
5 balance += 500
6 if balance > 1000:
7 status = 1
8 flag = True
9 limit = status * 5000
10 transaction = limit * 0.2
11 return flag
#### Example Question 1.1:
Is there information flow from (balance,4) to (transaction,10)? If so, provide a trace.
Analysis:
– Line 10: transaction reads limit; direct explicit flow (limit,9) -> (transaction,10).
– Line 9: limit reads status; direct explicit flow (status,7) -> (limit,9).
– Line 7: assignment executed only if line 6 condition true; control-dependent on line 6.
– Line 6: condition balance > 1000 reads balance; implicit flow (balance,5) -> (status,7).
– Line 5: balance updated from (balance,4); explicit flow (balance,4) -> (balance,5).
Output:
‘‘‘json
"InformationFlow": true,
"Trace": [
{ "from": ["balance", 4], "to": ["balance", 5], "type": "data" },
{ "from": ["balance", 5], "to": ["balance", 6, "use"], "type": "data" },
{ "from": ["balance", 6, "use"], "to": ["status", 7], "type": "control" },
{ "from": ["status", 7], "to": ["limit", 9], "type": "data" },
{ "from": ["limit", 9], "to": ["transaction", 10], "type": "data" }
#### Example Question 1.2:
Is there information flow from (flag,3) to (limit,9)? If so, provide a trace.
Analysis: No transitive chain (explicit or implicit) connects (flag,3) to (limit,9). Output:
‘‘‘json
"InformationFlow": false
#### Example Question 1.3:
Is there information flow from (status,2) to (transaction,10)? If so, provide a trace.
Analysis: One feasible explicit-flow chain is (status,2) -> (limit,9) -> (transaction,10). Output:
‘‘‘json
"InformationFlow": true,
"Trace": [
{ "from": ["status", 2], "to": ["limit", 9], "type": "data" },
{ "from": ["limit", 9], "to": ["transaction", 10], "type": "data" }
### Example 2 ‘‘‘python
1. val = 5
2. size = 3
3. arr = [0] * size
4. i = 0
5. j = 1
6. total = 0
7. if val > 2:
8. arr[j % size] = j
9. while i < size:
10. arr[j] += 1
11. score = arr[j+1]
12. total += score * 2
13. diff = total - score
14. j = (j + 1) % size
15. i += 1
16. last = arr[-1]
17. summary = diff + j
#### Example Question 2.1:
Is there information flow from (size,2) to (summary,17)? If so, provide a trace.
Analysis: Feasible chain (size,2) -> (size,9,"use") -> (j,14) -> (summary,17). Output:
‘‘‘json
"InformationFlow": true,
"Trace": [
{ "from": ["size", 2], "to": ["size", 9, "use"], "type": "data" },
{ "from": ["size", 9, "use"], "to": ["j", 14], "type": "control" },
{ "from": ["j", 14], "to": ["summary", 17], "type": "data" }
#### Example Question 2.2:
Is there information flow from (j,5) to (score,11)? If so, provide a trace.
Analysis: Direct explicit flow (j,5) -> (score,11) suffices. Output:
‘‘‘json
"InformationFlow": true,
"Trace": [
{ "from": ["j", 5], "to": ["score", 11], "type": "data" }
#### Example Question 2.3:
Is there information flow from (j,1) to (i,15)? If so, provide a trace.
Analysis: No explicit or implicit chain links (j,1) to (i,15). Output:
‘‘‘json
"InformationFlow": false
#### Example Question 2.4:
Is there information flow from (val,1) to (last,16)? If so, provide a trace.
Analysis: One feasible chain (val,1) -> (val,7,"use") -> (arr,8) -> (last,16) exists. Output:
‘‘‘json
"InformationFlow": true,
"Trace": [
{ "from": ["val", 1], "to": ["val", 7, "use"], "type": "data" },
{ "from": ["val", 7, "use"], "to": ["arr", 8], "type": "control" },
{ "from": ["arr", 8], "to": ["last", 16], "type": "data" }
[YOUR TURN] Below is your target snippet. “‘
<target code with line number>
Question: Is there information flow from (src_var, src_line) to (dst_var, dst_line) in function target_function_name? If so, provide a trace.
Output:
Prompt for Dependency Source Enumeration (Information Flow) You are a program analysis assistant. Please perform a static analysis of “information flow” on a given code snippet, treating each branch or loop condition as potentially taking any outcome, without using semantic or symbolic execution to prune paths. ## 1. Information Flow Definition An information flow from variable x to variable y, written x $\rightarrow$ y, exists whenever information stored in x is transferred to or used to derive information transferred to y. We denote each variable instance as (var,lineNumber), meaning the variable var is defined or updated at lineNumber. ### Types of Direct Flows Direct Explicit Flow A direct explicit flow occurs when the value written by one variable instance is directly used in computing the value of another, through operations like assignments. Formally, if (varA, lineA) writes a value and (varB, lineB) reads that value to compute its own, then there is a direct explicit flow from (varA, lineA) to (varB, lineB), written as: (varA,lineA) $\rightarrow$ (varB,lineB). Example 1:
1 x = 5
2 y = x + 1
(x,1) $\rightarrow$ (y,2) is a direct explicit flow. Example 2:
‘‘‘python
1 x = 5
2 if cond:
3 y = x + 1
(x,1) $\rightarrow$ (y,3) is still a direct explicit flow since static analysis assumes line 3 may execute.
Direct Implicit Flow
A direct implicit flow from (varA,lineA) to (varB,lineB) occurs when (varA,lineA)’s value directly influences whether lineB executes—that is, there is a direct control dependence.
Formally, if (varA,lineA) is read in a condition at lineC and that condition directly decides whether lineB executes, we write the direct implicit flow as (varA,lineA) $\rightarrow$ (varB,lineB). Example:
‘‘‘python
1 z = x + 5
2 if z > 1:
3 y = 10
(z,1) $\rightarrow$ (z,2)* $\rightarrow$ (y,3) is a direct implicit flow through the condition at line 2. The asterisk “*” marks that z is used at the condition and not re-defined there.
## 2. Output Format When asked “Which variable instances have information flow over (targetVar,targetLine)? List all such variables.”, respond: With variables present:
‘‘‘json
"InfomationFlowSources": [
["valA", lineA],
["valB", lineB],
With none present:
‘‘‘json
"InfomationFlowSources": false
A flow exists from (varA,lineA) to (varB,lineB) if there is any transitive chain of direct flows—explicit, implicit, or both—linking them. ## 3. Intraprocedural Data Dependence All analysis is confined to a single function; flows are not tracked across function boundaries. ## 4. Example Code Snippet ### Example 1 ‘‘‘python
1 def example_func():
2 status = 0
3 flag = False
4 balance = 1000
5 balance += 500
6 if balance > 1000:
7 status = 1
8 flag = True
9 limit = status * 5000
10 transaction = limit * 0.2
11 return flag
#### Example Question 1.1:
Which variable instances have information flow over (transaction,10)? List all such variables.
Analysis:
There is an explicit chain: (limit,9) $\rightarrow$ (transaction,10), then (status,7)/(status,2) $\rightarrow$ (limit,9). Line 7 is control-dependent on the condition at line 6, which reads balance. Thus flows propagate from (balance,4) and (balance,5).
Output:
‘‘‘json
"InfomationFlowSources": [
["limit", 9],
["status", 2],
["status", 7],
["balance", 4],
["balance", 5]
#### Example Question 1.2:
Which variable instances have information flow over (flag,3)? List all such variables.
Analysis: Line 3 simply initializes flag. Output:
‘‘‘json
"InfomationFlowSources": []
### Example 2 ‘‘‘python
1. val = 5
2. size = 3
3. arr = [0] * size
4. i = 0
5. j = 1
6. total = 0
7. if val > 2:
8. arr[j % size] = j
9. while i < size:
10. arr[j] += 1
11. score = arr[j+1]
12. total += score * 2
13. diff = total - score
14. j = (j + 1) % size
15. i += 1
16. last = arr[-1]
17. summary = diff + j
#### Example Question 2.1:
Which variable instances have information flow over (summary,17)? List all such variables.
Output:
‘‘‘json
"InfomationFlowSources": [
["diff", 13],
["j", 14],
["j", 5],
["size", 2],
["i", 4],
["i", 15]
#### Example Question 2.2:
Which variable instances have information flow over (score,11)? List all such variables.
Output:
‘‘‘json
"InfomationFlowSources": [
["arr", 10],
["arr", 8],
["arr", 3],
["j", 14],
["j", 5],
["size", 2],
["i", 4],
["i", 15],
["val", 1]
#### Example Question 2.4:
Which variable instances have information flow over (last,16)? List all such variables.
Output:
‘‘‘json
"InfomationFlowSources": [
["arr", 10],
["arr", 8],
["arr", 3],
["j", 14],
["j", 5],
["size", 2],
["i", 4],
["i", 15],
["val", 1]
[YOUR TURN] Below is your target snippet. “‘
<target code with line number>
Question: Which variable instances have information flow over (dst_var, dst_line) in function target_function_name? List all such variables.
Output:
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the paper’s goal of evaluating LLMs’ code reasoning capabilities through fundamental static analysis tasks. These claims align with the benchmark design, evaluation methodology, and experimental results presented in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are explicitly discussed in Section 8, including challenges with scaling and the potential for human annotation errors.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include any theoretical results or formal proofs; it focuses on benchmark design and empirical evaluation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We release the full benchmark, evaluation code, and experiment scripts to ensure reproducibility of results and verification of the main claims. They can be found in the released artifact.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release the full benchmark, evaluation code, and experiment scripts to ensure reproducibility of results and verification of the main claims. They can be found in the released artifact.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper involves inference-only experiments; details of model invocation and hyperparameter settings are provided in Section E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Our experiments use deterministic inference with temperature set to 0 to minimize randomness, following common practice in prior work [84, 56, 93].
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All experiments are conducted via third-party APIs; invocation and configuration details are provided in Section E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research adheres to all aspects of the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper includes a discussion of broader impacts, highlighting the positive impact of improving LLM reasoning skill assessment. We also discuss limitations.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper does not release any models or datasets with a high risk of misuse. The released benchmark consists of code snippets derived from public programming competition data, which poses minimal safety or misuse concerns.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We sampled data from two existing benchmarks, and their licenses are clearly acknowledged and discussed in Section A.1.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The paper introduces a new benchmark dataset, CoRe, which is thoroughly documented and released with metadata, task descriptions, annotation format, and usage instructions as part of the artifact.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve any crowdsourcing or human subject research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve human subjects or any form of crowdsourcing.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methodology and research contributions do not involve LLMs as an original or non-standard component.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.